### v2

# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

Abstract
While Graph Neural Networks (GNNs) have proven highly effective at modeling relational data, pairwise connections cannot fully capture multi-way relationships naturally present in complex real-world systems. In response to this, Topological Deep Learning (TDL) leverages more general combinatorial representations —such as simplicial or cellular complexes— to accommodate higher-order interactions. Existing TDL methods often extend GNNs through Higher-Order Message Passing (HOMP), but face critical scalability challenges due to (i) a combinatorial explosion of message-passing routes, and (ii) significant complexity overhead from the propagation mechanism.This work presents HOPSE (Higher-Order Positional and Structural Encoder), an alternative method to solve tasks involving higher-order interactions without message passing. Instead, HOPSE breaks apart arbitrary higher-order domains into it’s neighborhood relationships using a Hasse graph decompositions. This method shows that decoupling the representation learning of the topology of each neighborhood and the attributes results in a lower computational complexity, putting into question the need for HOMP.The experiments on molecular graph tasks and topological benchmarks show that HOPSE matches performance in traditional TDL datasets and outperforms HOMP methods in topological tasks, all while achieving up to $7\times$ speedups over HOMP-based models, opening a new path for scalable TDL.
1 Introduction
Graph neural networks (GNNs) (Zhou et al., 2020; Bronstein et al., 2021) have demonstrated impressive capabilities in modeling relational data structured as graphs. However, their reliance on edges to encode interactions inherently limits them to pairwise relationships, which may not fully capture the complexity of many real-world systems. For instance, social dynamics often involve group interactions, while physical phenomena like electrostatic interactions in proteins can span multiple atoms simultaneously.
To address such limitations, Topological Deep Learning (TDL) (Bodnar, 2023; Hajij et al., 2023) introduces a framework for modeling multi-way interactions using higher-order topological domains such as hypergraphs, simplicial, cell, and combinatorial complexes. Within TDL, standard Topological Neural Networks (TNNs) (Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; 2022; Papillon et al., 2023) extend the graph-based message passing paradigm through Higher-Order Message Passing (HOMP), enabling information flow across all of the dimensions of the considered combinatorial representation. [^1] The terms higher-order topological domain and combinatorial representation are used intercheangeably: please see Section A.1 for the formal definitions. This generalization has shown strong performance across diverse fields including social networks (Knoke and Yang, 2019), protein biology (Jha et al., 2022), physics-informed learning (Wei and Fink, 2024), and computer networks (Bernárdez et al., 2025; 2023) where interactions naturally extend beyond pairwise connections.
However, despite their theoretical advantages, TNNs face substantial challenges stemming from the HOMP paradigm. Unlike GNNs, TNNs’ ability to propagate information through the full hierarchy of a higher-order domain, where elements of different ranks (e.g., nodes, edges, faces, etc.) are interconnected, leads to an explosion in the number of possible sequences of message passing routes across these different levels (Telyatnikov et al., 2024; Montagna et al., 2024; Papamarkou et al., 2024). More specifically, for a combinatorial complex $\mathcal{T}$ with dimension —i.e. maximum rank— $R$, a total of $(R+1)!$ message passing routes arises just considering communication flows between different ranks, and that number skyrockets to $(R+1)!\cdot(R+1)^{R}$ when allowing same-rank connections (please see Appendix˜C for a detailed analysis).
On top of this vast architectural search space, TNNs operating under the HOMP paradigm are also known to be computationally demanding, message propagation often requiring significant processing resources even for moderately sized inputs (Papamarkou et al., 2024; Gurugubelli and Chepuri, 2024). This combination of combinatorial design complexity and high computational cost severely limits existing TNNs’ scalability and is not required for many tasks, even ones that require leveraging higher-order relationship. Appendix˜B contains theoretical motivation for this claim.
In response to these challenges, this work introduces the Higher-Order Positional and Structural Encoder (HOPSE), a novel and scalable framework that bypasses the HOMP paradigm altogether. While the message-passing preprocessing philosophy has been successfully applied within the simplicial domain (SaNN (Gurugubelli and Chepuri, 2024)), HOPSE does not rely on it altogether and achieves a critical generalization to arbitrary higher-order domains. By employing the Hasse graph decomposition as a methodological bridge, this framework enables the rich literature of powerful graph-based positional and structural encodings (PSEs) to be systematically and flexibly applied to any higher-order structure.
More specifically, HOPSE’s main features and contributions include:
  • Higher-order topological characterization: By decomposing a combinatorial complex into simpler graph ensembles, called Hasse graphs  (Hajij et al., 2023; Papillon et al., 2023), HOPSE efficiently uses established graph PSE techniques (e.g. (Cantürk et al., 2023)) to obtain characterizations for higher-order topological domains.
  • Scalability and computational efficiency: In contrast to HOMP networks, HOPSE complexity scales linearly with respect to the number of elements in the combinatorial representation (Section˜4.2). This is achieved by computing the encodings as a preprocessing step, significantly accelerating both training and inference (see Section˜5 for experimental details).
  • Extensive empirical validation: Comprehensive experiments on graph  (Morris et al., 2020) and topological  (Ballester et al., 2025) datasets validate the practical effectiveness of HOPSE. It is able to perform on-par to HOMP methods on graph tasks while showing stronger characterization capabilities required for topological tasks with a speed up of up to $7\times$ with respect to HOMP models (Section˜5).
Additionally, Appendix˜C provides, for the first time, a detailed mathematical analysis of the combinatorial explosion of message passing routes in higher-order topological domains.
2 Background
This section introduces key concepts underpinning TDL. For brevity and generality, the following discussion centers on combinatorial complexes (CCs), which subsume all of the discrete topological domains leveraged by TDL Hajij et al. (2023). Appendix A provides rigorous mathematical definitions of these other domains, such as simplicial and cellular complexes, along with an extended theoretical background.
Combinatorial complex.
A combinatorial complex $\mathcal{T}$} is a triple $(\mathcal{V},\mathcal{C},\textrm{rk})$ consisting of a set $\mathcal{V}$, a subset $\mathcal{C}$ of the powerset $\mathcal{P}(\mathcal{V})\backslash{\emptyset}$, and a rank function $\textrm{rk}:\mathcal{C}\rightarrow\mathbb{Z}_{\geq 0}$ with the following properties:
  1. 1.
    For all $v\in\mathcal{V},\{v\}\in\mathcal{C}$ and $\textrm{rk}(\{v\})=0$;
  2. 2.
    The function rk is order-preserving, i.e., if $\sigma,\tau\in\mathcal{C}$ satisfy $\sigma\subseteq\tau$, then $\textrm{rk}(\sigma)\leq$ $\textrm{rk}(\tau)$;
The elements of $\mathcal{V}$ represent the nodes, while the elements of $\mathcal{C}$ are called cells (i.e., groups of nodes). The cell $\sigma\in\mathcal{C}$ of rank $k:=\textrm{rk}(\sigma)$ is referred to as a $k$-cell. Lastly, the dimension of $\mathcal{T}$} is defined as the maximal rank among its cells: $\mathrm{dim}(\mathcal{C}):=\max_{\sigma\in\mathcal{C}}\textrm{rk}(\sigma)$.
Neighborhood relations: adjacencies and incidences.
Combinatorial complexes can be equipped with a notion of neighborhood among cells that induces topological structures. Formally, denoting by $\mathcal{P}(\cdot)$ the power set, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a complex $\mathcal{T}$} maps each cell $\sigma\in\mathcal{C}$ to a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}$. Traditional neighborhood functionsme include adjacencies ($\mathcal{A}_{t,s}$), connecting cells of the same rank $t$ —through adjacency relations w.r.t. cells of another rank ss—, and incidences ($\mathcal{I}_{s\to t}$), linking cells of different ranks from a source rank $s$ to a target rank $t$. Figure˜1 presents some instances, and Section˜A.2 provides formal definitions.
Figure 1: Neighborhood expansions of a complex. Given a complex $\mathcal{T}$} (left), three examples of strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}}$ are illustrated corresponding to 4 neighborhood functions: adjacency of nodes w.r.t. edges (${\mathcal{A}}_{0,1}$), incidence from nodes to edges (${\mathcal{I}}_{0\to 1}$), incidence from faces to nodes (${\mathcal{I}}_{2\to 1}$), and incidence from nodes to faces (${\mathcal{I}}_{0\to 2}$).
Strictly augmented Hasse graphs.
Given a complex $\mathcal{T}$}, each particular neighborhood function $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ (Papillon et al., 2024), defined as the directed graph whose nodes and edges are given, respectively, by
$\mathcal{C}_{{\mathcal{N}}}=\{\sigma\in{\mathcal{C}}\mid{\mathcal{N}}(\sigma)\neq\emptyset\}\quad\text{and}\quad\mathcal{E}_{{\mathcal{N}}}=\{(\tau,\sigma)\mid\sigma,\tau\in{\mathcal{C}}_{{\mathcal{N}}},\ \tau\in{\mathcal{N}}(\sigma)\}.$
Figure˜1 visually shows some examples of the graph expansions induced by these neighborhood functions. Notably, strictly comes from the fact that the set of nodes of $\mathcal{G}_{{\mathcal{N}}}$ is restricted to $\mathcal{C}_{{\mathcal{N}}}$, i.e., to those cells that have at least one neighbor according to $\mathcal{N}$}. In contrast to this, the original definition of augmented Hasse graph considers the full set of cells $\mathcal{C}$ (Hajij et al., 2023).
Rank-targeted neighborhood functions.
The rank-targeted neighborhood is defined as the collection of neighborhood functions whose target cells have the same fixed rank $r$: $\mathcal{N}^{\to r}:=\{\mathcal{N}\in\mathcal{N}_{S}\mid\forall\sigma\in\mathcal{C},\forall\tau\in\mathcal{N}(\sigma):\mathrm{rk}(\tau)=r\}$.
3 Related works
Graph characterization in GNNs.
In the graph domain, a range of positional and structural encodings (PSEs) have been proposed to enhance node representations and characterize graphs. Positional encodings typically capture a node’s location within the graph (e.g., via distances or coordinates), while structural encodings describe the node’s role or function in the topology (e.g., node degrees). Classical approaches of positional encodings include spectral methods based on Laplacian eigenvectors (LapPE) Dwivedi et al. (2023), while a common structural encoding is the random-walk embedding (RWSE) Dwivedi et al. (2022); Wang et al. (2022), which reflects multi-hop connectivity. These handcrafted features help characterize graphs with otherwise similar local neighborhoods. More recently, transformer-based GNNs have incorporated PSEs into attention mechanisms Dwivedi et al. (2021); Rampášek et al. (2022); Ma et al. (2023). Other works focus on decoupling the positional information from the structure and learn the embeddings instead of picking them manually (Dwivedi et al., 2021). Building upon this, the Graph Positional and Structural Encoder (GPSE) Cantürk et al. (2023) offers a unified and transferable solution by learning a shared latent space of PSEs through self-supervised pretraining. GPSE’s embeddings can be integrated into downstream GNNs, achieving performance on par with or exceeding task-specific encodings, and advancing the development of general-purpose graph models.
Efficient message passing methods
There are many different ways to achieve expressive and computationally efficient GNNs but the same cannot be said for TNNs. Various works in the GNN domain have approached message passing over subgraphs or partitions of the original graph (Zeng et al., 2021; Li et al., 2021). Additionally, other approaches focus on decoupling learning the features from the structure (Frasca et al., 2020; Doshi and Chepuri, 2022; Chen et al., 2020). In higher-order domains partitions are not trivial, so other methods have been applied to achieve efficiency.Simplicial Scattering Networks  (Madhu et al., 2024), leverage random walks over the intitial simplicial complex which give node structure-informed embeddings resulting in a parameter-free network. However, their experimental framework is mostly focused in node-level tasks. To the best of our knowledge Simplicial-aware Neural Networks (SaNN(Gurugubelli and Chepuri, 2024) is the only method that proposes efficient computations for higher-order networks in a learning scenario. They precompute a deterministic transformation over both the structure and the features of neighborhoods in a simplicial complex. Effectively, this results in a precomputation of message passing. However, it has several drawbacks. First, there is no clear indication of which transformation of the features and structure of a neighborhood should be used. Since this is a preprocessing step, this function cannot be learned. Second, this approach is limited to only simplicial complexes. HOPSE joins the ideas from (Frasca et al., 2020; Chen et al., 2020) and (Gurugubelli and Chepuri, 2024) to partially decouple learning the structure from the attributes. Additionally, HOPSE bypasses the theoretical limitations of (Gurugubelli and Chepuri, 2024) (and HOMP-based models), which is at most as expressive as SWL.
4 HOPSE: higher-order positional and structural encoder
Section˜4.1 outlines the proposed HOPSE pipelineand Section˜4.2 presents two practical realizations of the framework.
4.1 Methodology
This section describes HOPSE, a method for learning representations of (CCs) through the caracterization of it’s pieces. Characterizing a CC amounts to determining properties or invariants. These are shared among equivalnt CCs and allow distinguishing non-equivalent CCs from each toher, however similar they may be. The Hasse decomposition facilitates this characterization as a CC is determined by it’s Hasse decomposition. This decomposition results in a set of graphs that determines the initial object. For the theoretical motivation refer to Appendix˜B.
This work leverages the previous theoretical insights and the extensive research in graph characterization, for example through PSEs, to employ Hasse graphs as proxy objects to characterize a CC and model higher-order interactions. In the spirit of (Frasca et al., 2020), HOPSE partially decouples the structural from the functional representation. The structure is given by the Hasse decomposition while the functional representation is given by the attributes over the complex. Learning both structure and features in parallel circumvents the high cost of mixing features through neighborhood relationships, the pillar behind HOMP. The full pipeline is illustrated in Figure˜2, with components referenced using the corresponding labels (A–F) throughout this section.
Figure 2: HOPSE pipeline. (A) Considering the collection of neighborhood functions ${\mathcal{N}}_{C}=\{{\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{0\to 2}\}$, the input combinatorial complex $\mathcal{T}$ is decomposed into the corresponding Hasse graphs expansions $\{\mathcal{G}_{{\mathcal{N}}}\}_{{\mathcal{N}}_{C}}$. (B) PSEs are computed for each Hasse graph, $\mathbf{G}(\mathcal{G}_{\mathcal{N}_{i}})$, as defined in Equation˜1. (C) Higher-order PSEs are derived via the target rank-aware aggregation $\psi_{k}$ (see Equation˜2). (D) The PSEs $\mathbf{X}_{r,k}$ are transformed using $f_{r,k}(\cdot)$, and initial cell features $\mathbf{Z}_{r}$ are embedded via $\varepsilon_{r}(\cdot)$, following Equations˜3 and 4. (E) Rank-specific, structure-aware representations $\mathbf{H}_{r}$ are learned using $\Theta_{r}(\cdot)$, as described in Equation˜5. (F) The final task-specific readout $\varphi(\cdot)$ integrates all $\mathbf{H}_{r}$ representations (Equation˜6).
Extract positional and structural information.
Let ${\mathcal{N}}_{S}$ be a collection of neighborhood functions. For each neighborhood function ${\mathcal{N}}\in{\mathcal{N}}_{S}$, the Hasse graph decomposition of $\mathcal{T}$} enables the treatment of every ${\mathcal{G}}_{{\mathcal{N}}}$ as an independent component (see Figure˜2A), each representing a unique set of connectivity relationships. $\mathbf{G}$ is a set of Positional and Structural Encoders (PSEs) such that
$\mathbf{G}\colon\{1,\dots,K\}\times\mathcal{G}_{S}\longrightarrow\mathbb{R}^{N\times D_{k}},\quad\left(k,\mathcal{G}_{\mathcal{N}}\right)\longmapsto g_{k}\left(\mathcal{G}_{\mathcal{N}}\right).$ (1)
Here $g_{k}(\cdot):\mathcal{G}_{S}\to\mathbb{R}^{|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}\times D_{k}}$ denotes a particular PSE, with $|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}$ being the number of nodes in the Hasse graph constructed from the neighborhood function $\mathcal{N}$} corresponding to the target cells (illustrated in Figure˜2B). Note that each $g_{k}(\cdot)$ maps to a different $D_{k}$, allowing PSEs to have different dimensions.
Aggregate positional and structural encodings across neighborhoods.
The function $\psi_{k}(\cdot)$ groups encodings of cells with the same target rank $r$ to produce a embedding that characterizes each cell and for each $k\in\{1,\dots,K\}$ (see Figure˜2C). The embeddings for cells of rank $r$ are computed as:
$\mathbf{X}_{r,k}=\psi_{k}(\{\!\!\{g_{k}(\mathcal{G}_{\mathcal{N}})\mid{\mathcal{N}}\in{\mathcal{N}}^{\to r}\}\!\!\}),$ (2)
where $\mathbf{X}_{r,k}\in\mathbb{R}^{N_{r}\times D_{k^{\prime}}}$, with $N_{r}$ denoting the number of $r$-cells and $D_{k^{\prime}}$ representing the output feature dimension for attribute $k$. The functions $\psi_{k}(\cdot)$ can also be interpreted as rank-level aggregators, where the aggregation operation can be any standard reduction method over a multiset  [^2] Common choices include concatenation, sum, average, or even learnable transformations such as MLPs.. The main advantage of using non-learnable aggregation functions for $\psi_{k}$ is to precompute the embeddings $\mathbf{X}_{r,k}$ during dataset preprocessing.
Learning structure and attributes.
In the HOPSE framework, each $r$-cell is associated with two complementary sources of information: (i) neighborhood-aware PSEs $\mathbf{X}_{r,k}$ , and (ii) initial features $\mathbf{Z}_{r}$ provided by the dataset or derived during the construction of the higher-order structure. Both types of features are projected in parallel into latent spaces through sets of learnable functions (see Figure˜2D):
$\hat{\mathbf{X}}_{r,k}$ $=f_{r,k}(\mathbf{X}_{r,k}),$ (3)
$\hat{\mathbf{Z}}_{r}$ $=\varepsilon_{r}(\mathbf{Z}_{r}),$ (4)
where $f_{r,k}(\cdot)$ is applied to each structural feature matrix $\mathbf{X}_{r,k}$ of each $r$-cell, and $\varepsilon_{r}(\cdot)$ to their initial features.
Learning neighborhood-aware representations.
Next, to integrate the structural information with the cells’ initial attributes, $\Theta_{r}(\cdot)$ aggregates their latent representations (see Figure˜2E) to obtain the final neighborhood-aware representations $\mathbf{H}_{r}\in\mathbb{R}^{N_{r}\times D_{r}}$:
$\mathbf{H}_{r}=\Theta_{r}(\hat{\mathbf{Z}}_{r},\hat{\mathbf{X}}_{r,1},\dots,\hat{\mathbf{X}}_{r,K}),$ (5)
Learning the downstream task.
Since the structural information is now contained in the cell features, they can be treated independently. For example, for complex-level tasks where the network needs an output for the whole complex, the global transformation $\varphi(\cdot)$ combines the aggregated embeddings from all ranks (see Figure˜2F). The overall output is
$\text{HOPSE}({\mathcal{T}})=\varphi\Bigl(\mathbf{H}_{0},\ldots,\mathbf{H}_{R}\Bigr).$ (6)
4.2 Realization of the framework
The practical implementation of the proposed framework requires defining a set of $T$ neighborhood functions of interest ${\mathcal{N}}_{S}=\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$ (discussed in Section˜5), a set of encoding functions $\mathbf{G}$ for deriving structural and positional features, neighborhood aggregation functions $\psi_{k}(\cdot)$, rank-specific processing functions $f_{r,k}(\cdot)$, and a structure mixing function $\Theta_{r}(\cdot)$. In particular, this work adopts $\psi_{k}(\cdot)=\mathrm{CONCAT}(\cdot)$ for aggregation, enabling each positional encoding to be handled independently. The functions $\epsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$ and $\Theta_{r}(\cdot)$ are modeled as multilayer perceptrons (MLPs). The choice of functions $\mathbf{G}$ defines two instantiations of the framework, HOPSE-M and HOPSE-G, offering a trade-off between representational power  [^3] Since HOPSE-G uses message passing to generate the embeddings of the Hasse graphs it is not completely message passing free.. and computation cost. Additional implementation details are provided in Appendix˜D.
HOPSE-M.
This realization incorporates manually selected positional and structural encodings commonly used in graph learning. Specifically, $\mathbf{G}=\{g_{1},g_{2},g_{3},g_{4}\}$ includes: (i) eigenvector-based encodings, (ii) electrostatic potential-based encodings, (iii) random-walk-based encodings, and (iv) heat kernel signatures. Section˜F.2 contains an in depth description of the choice of PSEs.
HOPSE-G.
This variant leverages the pretrained GPSE network (Cantürk et al., 2023) to encode structural and positional information. In this realization, $\mathbf{G}=\{g_{1}\}$ with $g_{1}=\textsc{GPSE}(\cdot)$. Unlike manual encodings that rely on computationally intensive operations such as eigendecomposition, GPSE produces latent descriptors using standard message passing. Domain specific tasks can thus benefit from models pretrained in that domain.
Choice of ${\mathcal{N}}_{S}$.
The choice of the set of neighborhood functions ${\mathcal{N}}_{S}$ used by HOPSE is a critical design choice in the HOPSE framework. Following the theoretical motivation (Appendix˜B) extensive ablation studies show the empirical performance (Appendix˜H) and runtime behavior (Appendix˜I) of the models.
Computational complexity.
The HOPSE framework is efficient. No training is required up to and including Equation˜2, allowing the matrices $\mathbf{X}_{r,k}$ to be fully precomputed in a preprocessing phase. By decoupling the learning procedure of features and structure, HOPSE has complexity that scales linearly with the number of cells.
5 Experiments
This section empirically evaluates the HOPSE framework against state-of-the-art TNNs. Both instantiations of the framework, HOPSE-M and HOPSE-G (see Section 4.2), are assessed on graph molecular datasets and topological tasks. This work adopts the standardized experimental protocol introduced in TopoBench (Telyatnikov et al., 2024) to ensure fair and reproducible comparisons through a comprehensive optimization and evaluation suite. The experimental setup is carefully designed to systematically address the following research questions:
  1. Q1.
    Is HOMP required for standard TDL evaluation benchmarks ?
  2. Q2.
    Is HOMP a hard requirement for tasks related to topological properties?
  3. Q3.
    Does the lower computational complexity result in a reduced training and inference time ?
5.1 Evaluation setup
Datasets.
First, the framework is evaluated on five standard benchmarking datasets commonly used in the TDL literature (Telyatnikov et al., 2024), comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^4] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluation.MUTAG, PROTEINS, NCI1, and NCI109 (Morris et al., 2020), and the regression is ZINC-12K (Irwin et al., 2012). Following standard practice in TDL, structural lifting is applied to transform these graph datasets into higher-order domains—specifically, cycle-based lifting for the cell domain and clique complex lifting for the simplicial domain. Next, the MANTRA set of datasets (Ballester et al., 2025) represents purely topological datasets and tasks. It comprises over 43,000 triangulations of 2-manifolds. Among the multiple classification tasks that MANTRA datasets offer, this analysis focuses on the following three: predicting the homeomorphism type (NAME), predicting the orientability of a triangulated surface (ORIENT), and assigning a numerical label to the first three Betti numbers $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. For further discussion of the chosen benchmarks refer to Section˜F.2.
Models.
The benchmark includes several models from the simplicial and cellular domains. In the simplicial domain: SCN (Yang et al., 2022), SCCNN (Yang and Isufi, 2023), and GCCN (Papillon et al., 2024). In the cellular domain: CCCN (Hajij et al., 2020), CWN (Bodnar et al., 2021), and GCCN. This work also provides a re-implementation of the SaNN architecture Gurugubelli and Chepuri (2024), as its source code was unavailable. Additionally, the first baseline results for the MANTRA family (Ballester et al., 2025) under the TopoBench protocol. Full details on the hyperparameter search and selected configurations are provided in Appendix F.3.
Training and evaluation.
Publicly available results are adopted from Telyatnikov et al. (2024) and Papillon et al. (2024) where applicable for graph datasets. While for the rest, the TopoBench evaluation protocol was applied, which consists of stratified splits with 50%/25%/25% train/validation/test ratios, except on ZINC-12K, which uses predefined splits (Irwin et al., 2012). The performance for each dataset is the result of the test performance of the highest validation, all averaged over 5 seeds. Across all tables, results are shown as mean $\pm$ standard deviation. The best results per domain are highlighted in bold, the best overall in red and those within one standard deviation of the best are in blue. The evaluation metrics are classification accuracy for MUTAG, PROTEINS, NCI1, NCI109, mean absolute error (MAE) for ZINC-12K, and F1 score for MANTRA. This standardized protocol ensures all methods are evaluated on an even playing field, which may lead to results that differ from some prior publications. Please refer to Section˜F.2 for further details. Additionally, Appendix˜K contains the full parameter budget of the best models.
5.2 Experimental Results
Table 1: Performance (ACC,MAE) and Total training time in seconds (T) on Graph Tasks.
Model Dataset MUTAG PROTEINS NCI1 NCI109 ZINC-12K ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) ACC ($\uparrow$) T ($\downarrow$) MAE ($\downarrow$) T ($\downarrow$) Graph GCN $69.79\pm 0.06$.80) $4\pm 1.$ $75.70\pm 0.02$.14) $8\pm 2.$ $72.86\pm 0.00$.69) $53\pm 20.$ $72.20\pm 0.01$.22) $37\pm 9.$ $0.62\pm 0.00$.01) $146\pm 28.$ GCN+PSEs $84.26\pm 0.03$.56) $3\pm 1.$ $75.77\pm 0.02$.98) $23\pm 5.$ $74.88\pm 0.00$.54) $106\pm 47.$ $74.70\pm 0.00$.79) $122\pm 39.$ $0.43\pm 0.00$.03) $903\pm 40.$ GAT $2.77\pm 0.02.77$) $4\pm 1.$ $76.34\pm 0.01$.66) $8\pm 2.$ $75.00\pm 0.01$.00) $57\pm 17.$ $73.80\pm 0.00$.73) $56\pm 9.$ $0.61\pm 0.00$.01) $171\pm 64.$ GAT+PSEs $80.43\pm 0.05$.08) $5\pm 1.$ $75.48\pm 0.03$.44) $36\pm 15.$ $77.86\pm 0.01$.02) $210\pm 28.$ $75.57\pm 0.00$.60) $212\pm 89.$ $0.39\pm 0.00$.02) $708\pm 83.$ Simplicial SCCNN $76.17\pm 0.06$.63) $14\pm 3.$ $74.19\pm 0.02$.86) $53\pm 11.$ $76.60\pm 0.01$.75) $307\pm 83.$ $77.12\pm 0.01$.07) $354\pm 106.$ $0.36\pm 0.00$.02) $2060\pm 408.$ GCCN $85.96\pm 0.04$.66) $61\pm 7.$ $74.91\pm 0.02$.51) $66\pm 21.$ $76.85\pm 0.01$.50) $904\pm 180.$ $75.76\pm 0.01$.28) $538\pm 39.$ $0.35\pm 0.00$.01) $3603\pm 475.$ HOPSE-M (Our) $86.38\pm 0.06$.31) $10\pm 3.$ $75.13\pm 0.03$.66) $44\pm 17.$ $77.35\pm 0.00$.81) $201\pm 88.$ $76.34\pm 0.01$.40) $179\pm 79.$ $0.39\pm 0.00$.02) $1332\pm 721.$ HOPSE-G (Our) $85.53\pm 0.02$.77) $5\pm 1.$ $76.27\pm 0.02$.10) $33\pm 8.$ $77.70\pm 0.01$.40) $130\pm 22.$ $76.36\pm 0.01$.75) $107\pm 28.$ $0.43\pm 0.00$.01) $478\pm 107.$ Cell GCCN $86.38\pm 0.06$.49) $61\pm 18.$ $4.41\pm 0.01.77$) $59\pm 18.$ $78.23\pm 0.01$.47) $523\pm 119.$ $77.19\pm 0.00$.21) $386\pm 76.$ $0.19\pm 0.00$.01) $3301\pm 440.$ HOPSE-M (Our) $88.09\pm 0.05$.55) $7\pm 3.$ $76.99\pm 0.02$.63) $35\pm 17.$ $78.09\pm 0.01$.44) $156\pm 90.$ $77.27\pm 0.01$.01) $147\pm 81.$ $0.24\pm 0.00$.01) $760\pm 533.$ HOPSE-G (Our) $88.09\pm 0.01$.90) $7\pm 1.$ $76.42\pm 0.02$.19) $32\pm 5.$ $79.07\pm 0.00$.65) $153\pm 8.$ $77.81\pm 0.01$.26) $104\pm 38.$ $0.21\pm 0.00$.01) $1230\pm 527.$
Table 2: Performance (F1) and Total training time in seconds (T) on Topological Tasks.
Model Task NAME ORIENT $\beta_{1}$ $\beta_{2}$ F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) F1 ($\uparrow$) T ($\downarrow$) Graph GIN $76.14\pm 0.00$.14) $3568\pm 44.$ $56.28\pm 0.00$.45) $3536\pm 46.$ $88.13\pm 0.00$.00) $3674\pm 176.$ $0.93\pm 0.01$.21) $3674\pm 176.$ GIN+PSEs $94.05\pm 0.00$.15) $2174\pm 306.$ $82.53\pm 0.00$.78) $2128\pm 247.$ $90.28\pm 0.00$.05) $3011\pm 562.$ $68.43\pm 0.00$.59) $3011\pm 562.$ Simplicial SCCNN $95.08\pm 0.00$.56) $5521\pm 319.$ $86.29\pm 0.01$.23) $3536\pm 46.$ $90.20\pm 0.00$.20) $6720\pm 10.$ $65.82\pm 0.02$.70) $6720\pm 10.$ GCCN $86.76\pm 0.01$.27) $5898\pm 1961.$ $76.60\pm 0.01$.67) $4564\pm 631.$ $84.20\pm 0.04$.80) $7689\pm 2522.$ $41.82\pm 0.20$.19) $7689\pm 2522.$ HOPSE-M (Our) $96.27\pm 0.00$.28) $1530\pm 1366.$ $89.24\pm 0.00$.22) $1886\pm 1449.$ $91.02\pm 0.01$.43) $3179\pm 1756.$ $82.29\pm 0.00$.46) $3179\pm 1756.$ HOPSE-G (Our) $81.75\pm 0.01$.26) $3324\pm 1113.$ $62.17\pm 0.00$.98) $2795\pm 112.$ $88.28\pm 0.00$.08) $3876\pm 1141.$ $35.37\pm 0.02$.25) $3876\pm 1141.$
HOMP is not a hard requirement for performance.
Table˜1 and Table˜2 present the performance in graph and topological tasks respectively. The performance metric is the left most column for each dataset. In both scenarions, HOPSE-M and HOPSE-G models demonstrate consistently strong performance. This results illustrates the both Q1. and Q2. are negative, HOMP is not a requirement to capture higher-order relationships in traditional TDL benchmarks or in tasks related to topological properties. On standard graph classification benchmarks used in TDL, both HOPSE variants match or outperform most baselines, with HOPSE-M ranking among the top models and achieving state-of-the-art results in several cases (PROTEINS, NCI1, NCI109). While the performance gap on ZINC highlights the advantage of message-passing in fine-grained molecular regression tasks, HOPSE still remains competitive without relying embbedings generated with HOMP. The most significant gains are observed on the pure topological datasets. HOPSE-M delivers the strongest results on $\beta_{1}$ and $\beta_{2}$, outperforming all baselines, including models with explicit simplicial message passing. This suggests that HOPSE can effectively encode and leverage topological structure. It is also noteworthy that Table˜2 shows that GIN+PSEs always outperforms GCCNs and in $\beta_{1},\beta_{2}$ it outperforms SCCN. This results put into question if HOMP is usefull at all for topological tasks. For performance over all models refer to Appendix˜G, for ablations over different neighborhoods to Appendix˜H and for ablations over PSEs to Appendix˜J.
Speed does not need to compromise performance.
Tables˜2 and 1 show the total runtime of training in the rightmost column of each dataset.  [^5] This does not included preprocessing times, see Section I.3 for more in depth explanation.. In the category of models operating over simplicial or cell complexes, a HOPSE realization always achieves the fastest overall runtime. On graph datasets, the speedup can reach up to $7\times$—for example, HOPSE-G runs in 130 seconds compared to 904 seconds for GCCN on NCI1, while also delivering better performance. HOPSE’s efficiency becomes apparant as tasks scale in size and complexity. In NCI1, NCI109 and ZINC it is from $3\times$ to $7\times$ faster than other models with comparable or better performance. In topological tasks this is exacerbated, where HOPSE-M is up to $1.5\times$ more efficienct than GIN+PSEs while outperforming all other models. These results show that the theoretical complexity does result in empirically reduced training time, even when comparing to the case of GNNs. For runtimes over all models refer to Appendix˜G, for ablations over different neighborhoods to Appendix˜I and for model sizes to Appendix˜K.
6 Concluding remarks
Conclusions.
This work introduces HOPSE, a scalable and compute efficient framework designed to extract and leverage structural information of the Hasse decomposition to characterize arbitrary combinatorial representations. HOPSE performs competitively with HOMP without performing any type of message passing. Moreover, model scaling is linear with dataset size due parallel learning of structure and attributes. These claims are supported by experiments on molecular graph benchmarks, as well as and topological tasks, where both instantiations—HOPSE-M and HOPSE-G—match or surpass state-of-the-art performance while achieving up to $7\times$ speed-ups over HOMP-based models. By providing a characterization of a complex object via it’s constituting objects, Hasse graphs, HOPSE provides a new method for solving topological tasks. More importantly, in puts into question the necessity of HOMP. The cost of operating over augmented neighborhood structures blows up complexity without improving the model’s ability to learn the topology of the data. In a novel comparison, the ablations over GNNs with PSEs agains HOMP show that in several cases there is no need for the latter.
Nevertheless, this work not only enables broader applicability of TDL but also facilitates the integration of higher-order interactions into other areas of deep learning—mirroring the widespread adoption of graph-based methods (Zhou et al., 2023; Wang et al., 2025; Han et al., 2022; 2023).
Limitations & future work.
A deeper investigation of HOPSE encodings—along with a comprehensive comparison against alternative methods across diverse domains—lies beyond the scope of this work but remains an important and intriguing direction for future research. Other promising avenues include exploring graph-based strategies for generating higher-order positional and structural encodings, or better methods for characterizing Hasse graphs like the ones based on homomorphism counts. Additionally, this work does not address whether directly manufacturing encodings on a combinatorial complex provides a better characterization the joint characterization using it’s Hasse decomposition.Finally, and from a more practical standpoint, HOPSE-G realization of the framework leverages a pretrained model to extract latent structural and positional features, but it would be interesting to explore if a joint end-to-end training may further enhance performance –even at the cost of additional computational overhead.
Appendix A Extended background
A.1 Topological domains
This section draws primarily from Appendix A of Telyatnikov et al. [2024]. It begins with the fundamental concept of featured graphs, which forms the basis for understanding more intricate structures. Then it proceeds to explore higher-order domains — including simplicial complexes, cell complexes, and combinatorial complexes — each providing distinct capabilities for representing various types of relationships and hierarchies within data.
Definition A.1.
Let $\mathcal{G}=(V,E)$ be a graph, with node set $V$ and edge set $E$. A featured graph is a tuple $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$, where $F_{V}:V\to\mathbb{R}^{d_{v}}$ is a function that maps each node to a feature vector in $\mathbb{R}^{d_{v}}$ and $F_{E}:E\to\mathbb{R}^{d_{e}}$ is a function that maps each edge to a feature vector in $\mathbb{R}^{d_{e}}$.
A topological domain is a graph generalization that captures pairwise and higher-order relationships between entities [Bick et al., 2023, Battiston et al., 2021]. When working with topological domains, two key properties come into play: set-type relations and hierarchical structures represented by rank functions [Hajij et al., 2023, Papillon et al., 2023].
Definition A.2 (Set-type relation).
A relation in a topological domain is called a set-type relation if another relation in the domain does not imply its existence.
Definition A.3 (Rank function).
A rank function on a higher-order domain $\mathcal{X}$ is an order-preserving function $rk\colon\mathcal{X}\to\mathbb{Z}_{\geq 0}$ such that $x\subseteq y$ implies $rk(x)\leq rk(y)$ for all $x,y\in\mathcal{X}$.
Set-type relations emphasize the independence of connections within a domain, allowing for flexible representation of complex interactions. In contrast, rank functions introduce a hierarchical (also called part-whole) organization that facilitates the representation and analysis of nested relationships.
Simplicial complexes.
Simplicial complexes extend graphs by incorporating hierarchical part-whole relationships through the multi-scale construction of cells. In this structure, nodes correspond to rank 0-cells, which can be combined to form edges (rank $1$-cells). Edges can then be grouped to form faces (rank 2 cells), and faces can be combined to create volumes (rank $3$-cells), continuing in this manner. Consequently, the faces of a simplicial complex are triangles, volumes are tetrahedra, and higher-dimensional cells follow the same pattern. A key feature of simplicial complexes is their strict hierarchical structure, where each $k$-dimensional simplex is composed of (k1)(k-1)-dimensional simplices, reinforcing a strong sense of hierarchy across all levels.
Definition A.4 (Simplicial Complex).
A simplicial complex (SC) in a non-empty set $S$ is a pair $SC=(S,\mathcal{X})$, where $\mathcal{X}\subset\mathcal{P}(S)\setminus\{\emptyset\}$ satisfies: if $x\in SC$ and $y\subseteq x$, then $y\in SC$. The elements of $\mathcal{X}$ are called simplices.
Example 1 (3D Surface Meshes).
3D models of objects, such as those used in computer graphics or for representing anatomical structures, are often constructed using triangular meshes. These meshes naturally form simplicial complexes, where the vertices of the triangles are 0-simplices, the edges are 1-simplices, and the triangular faces themselves are 2-simplices.
Cell complexes
Cell complexes provide a hierarchical interior-to-boundary structure, offering clear topological and geometric interpretations, but they are not based on set-type relations. Unlike simplicial complexes, cell complexes are not limited to simplexes; faces can involve more than three nodes, allowing for a more flexible representation. This increased flexibility grants cell complexes greater expressivity compared to simplicial complexes Bodnar et al. [2021], Bodnar [2023].
Definition A.5 (Cell complex).
A regular cell complex is a topological space $S$ partitioned into subspaces (cells) $\{x_{\alpha}\}_{\alpha\in P_{S}}$, where $P_{S}$ is an index set, satisfying:
  1. 1.
    $S=\cup_{\alpha\in P_{S}}\text{int}(x_{\alpha})$, where $\text{int}(x)$ denotes the interior of cell $x$.
  2. 2.
    For each $\alpha\in P_{S}$, there exists a homeomorphism $\psi_{\alpha}$ (attaching map) from $x_{\alpha}$ to $\mathbb{R}^{n_{\alpha}}$ for some $n_{\alpha}\in\mathbb{N}$. The integer $n_{\alpha}$ is the dimension of cell $x_{\alpha}$.
  3. 3.
    For each cell $x_{\alpha}$, the boundary $\partial x_{\alpha}$ is a union of finitely many cells of strictly lower dimension.
Example 2 (Molecular structures.).
Molecules admit natural representations as cell complexes by considering atoms as nodes (i.e., cells of rank zero), bonds as edges (i.e., cells of rank one), and rings as faces (i.e., cells of rank two).
Combinatorial complexes.
Combinatorial complexes combine hierarchical structure with set-type relations, enabling a flexible yet comprehensive representation of higher-order networks.
Definition A.6 (Combinatorial complex).
A combinatorial complex (CC) is a triple $(\mathcal{V},\mathcal{C},{\textnormal{k}})$ consisting of a set $\mathcal{V}$, a subset $\mathcal{C}\subset\mathcal{P}(\mathcal{V})\setminus\{\emptyset\}$, and a function $rk\colon\mathcal{C}\to\mathbb{Z}_{\geq 0}$ satisfying:
  1. 1.
    For all $v\in\mathcal{V}$, $\{v\}\in\mathcal{C}$ and $rk(v)=0$.
  2. 2.
    The function $rk$ is order-preserving: if $x,y\in\mathcal{C}$ with $x\subseteq y$, then $rk(x)\leq rk(y)$.
Example 3 (Geospatial structures.).
Geospatial data, comprised of grid points (0-cells), road polylines (1-cells), and census tract polygons (2-cells), can be effectively represented using combinatorial complexes. A visual example is provided in Figure 2 (Right) of Battiloro et al. [2025].
Featured topological domains.
A featured graph is a graph whose nodes or edges are equipped with feature functions [Sanchez-Lengeling et al., 2021]. Similarly, each topological element (e.g., simplex or cell) can carry feature vectors to represent some information.
Definition A.7 (Featured topological domain).
A featured topological domain is a pair $(\mathcal{X},F)$, where $\mathcal{X}$ is a topological domain and $F=\{F_{i}\}_{i\geq 0}$ is a collection of feature functions. Each function $F_{i}$ maps the $i$-dimensional elements of $\mathcal{X}$, denoted $\mathcal{X}_{i}$, to a feature space $\mathbb{R}^{k_{i}}$:
$F_{i}\colon\mathcal{X}_{i}\to\mathbb{R}^{k_{i}}.$
A.2 Neighborhood functions and augmented Hasse graphs
Discrete topological domains are equipped with a notion of neighborhood among cells that confers on them a topological structure; this section revisits the main neighborhood-related concepts in the general combinatorial complex case.
Definition A.8 (Neighborhood function).
Given a combinatorial complex ${\mathcal{T}}=(\mathcal{V},\mathcal{C},{\textnormal{k}})$, a neighborhood function $\mathcal{N}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{C})$ on a combinatorial complex is a function that assigns to each cell $\sigma\in\mathcal{C}$ a collection of neighbor cells $\mathcal{N}(\sigma)\subset\mathcal{C}\cup\emptyset$.
Each neighborhood function can be used to create a Strictly Augmented Hasse graph.
Definition A.9 (Strictly augmented Hasse graph).
Each particular neighborhood function $\mathcal{N}$ induces a strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ on $\mathcal{T}$ [Papillon et al., 2024, Hajij et al., 2023], defined as the directed graph whose nodes and edges are given, respectively, by $\mathcal{C}_{\mathcal{N}}=\{\sigma\in\mathcal{C}\,|\,\mathcal{N}(\sigma)\neq\emptyset\}$ and $\mathcal{E}_{\mathcal{N}}=\{(\tau,\sigma)\,|\,\tau\in\mathcal{N}(\sigma)\}$.
Examples of neighborhood functions are incidences (connecting cells with different ranks) and adjacencies (connecting cells with the same rank), although other neighborhood functions can be defined for specific tasks [Battiloro et al., 2025]. More in detail, given a combinatorial complex $\mathcal{T}$}:
Definition A.10.
(Incidences) The incidence neighborhood of a $t$-cell $\sigma\in\mathcal{T}$ with respect to a rank $s$ is defined as
[…]
A $s$-cell $\tau\in\mathcal{T}$ is said to be co-incident of $\sigma$ if it contains $\sigma$ and $s>t$; analogously, a $s$-cell $\tau\in\mathcal{T}$ is incident of $\sigma$ if it is is contained by $\sigma$ and $s<t$.
Incidences in turn induce adjacency neighborhoods of cells of the same rank as follows:
Definition A.11 (Adjacencies).
The adjacency neighborhood of a $t$-cell $\sigma\in\mathcal{T}$ with respect to a rank $s$ is defined as
$\begin{split}\mathcal{A}_{s,t}(\sigma)=\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=t,\mathcal{I}_{s\to t}(\sigma)\cap\mathcal{I}_{s\to t}(\tau)\neq\emptyset\}\end{split}$ (8)
In this case, two $t$-cells $\sigma$ and $\tau$ are said to be adjacent if they are both contained in a $s$-cell $\delta\in\mathcal{T}$.
A.3 Morphisms between CCs
Before defining isomorphism between CCs, the definition of homomorphism between CCs is needed
Definition A.12 (CC-Homomorphism (Definition 10 Hajij et al. [2023]).
A homomorphism from a $CC(S_{1},\mathcal{X}_{1},\mathrm{rk}_{1})$ to another $CC(S_{2},\mathcal{X}_{2},\mathrm{rk}_{2})$ is a function $f:\mathcal{X}_{1}\to\mathcal{X}_{2}$ that sasfies the following conditions:
  1. 1.
    If $x,y\in\mathcal{X}_{1}$ satisfy $x\subseteq y$, then $f(x)\subseteq f(y)$
  2. 2.
    If $x\in\mathcal{X}_{1}$, then $\mathrm{rk}_{1}(x)\geq\mathrm{rk}_{2}(f(x))$.
However, it is not enough to work in the attributed setting. A CC with attributes assigned to cells is
Definition A.13 (Attributed CC [Papillon et al., 2024]).
An attributed combinatorial complex $({\mathcal{C}},\ell)$ is a CC denoted $\mathcal{C}$} equipped with a cell coloring function $\ell:{\mathcal{C}}\to\Sigma$ where $\Sigma$ is an arbitrary normed space. For a cell $\sigma$ that is part of $\mathcal{C}$}, $\ell(\sigma)$ denotes the label or color of $\sigma$.
Now note that there also exist homomorphisms induced by the neighborhood functions that can be applied to the CCs. The definition of such is
Definition A.14 (CC-Homomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ [Papillon et al., 2024]).
A homomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ with a neighborhood function $\mathcal{N}_{1}$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ with a neighborhood function $\mathcal{N}_{2}$ is a function $f:\mathcal{C}_{1}\to\mathcal{C}_{2}$ that satisfies the following condition: If $\sigma,\tau\in\mathcal{C}_{1}$ and $\tau\in\mathcal{N}_{1}(\sigma)$ then $f(\tau)\in\mathcal{N}_{2}(f(\sigma))$.
This definition extends to attributed CCs with the following condition: if $\sigma,\tau\in\mathcal{C}_{1}$ have the same label then their image in $\mathcal{C}_{2}$ also has the same label, i.e for $\ell:\mathcal{C}_{1}\to\Sigma$ and $\ell^{\prime}:\mathcal{C}_{2}\to\Sigma$, if $\ell(\sigma)=\ell(\tau)$ then $f(\ell^{\prime}(\sigma))=f(\ell^{\prime}(\tau))$. The neighborhood functions also induce Hasse graphs. In fact, homomorphisms between CCs induced by neighborhood functions are one-to-one to homomorphisms of the Hasse graphs induced by the same neighborhood functions.
Proposition A.15 (Proposition B.4 in [Papillon et al., 2024]).
For every CC-Homomorphism $f$ from $\mathcal{C}_{1}$ to $\mathcal{C}_{2}$ induced by the neighborhood functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, with domains $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, there exists a unique homomorphism between the strictly augmented Hasse graphs $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$.
The generality of combinatorial complexes complicates their direct isomorphic comparison. Interestingly, they can be decomposed and evaluated on their constituting pieces, Hasse graphs. Thus, isomorphisms can be induced by neighborhood functions, which have a one-to-one correspondance with Hasse graphs, which give an equivalence relationship between CCs. This type of equivalence relationship is definited as
Definition A.16 (CC-Isomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ [Papillon et al., 2024]).
A isomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ with a neighborhood function $\mathcal{N}_{1}$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ with a neighborhood function $\mathcal{N}_{2}$ is an invertible CC-Homomorphism induced by $(\mathcal{N}_{1},\mathcal{N}_{2})$ whose inverse is a CC-Isomorphism induced by $(\mathcal{N}_{2},\mathcal{N}_{1})$. In the labeled CC case, then the map also preserves labels.
Definition A.17 (CC-Isomorphism).
A isomorphism from a $CC$ denoted $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ to another $CC$ denoted $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$ is a bijection $f$ that fulfils:
  • $\forall\sigma,\tau\in\mathcal{C}_{1}$, if $\sigma\subseteq\tau$ then $f(\sigma)\subseteq f(\tau)$,
  • $\forall\sigma\in\mathcal{C}_{1}$ it holds that $\mathrm{rk}_{1}(\sigma)=\mathrm{rk}_{2}(f(\sigma))$.
If there exists such an $f$, the CCs are deemed isomorphic, denoted $CC_{1}\simeq CC_{2}$.
A.4 Characterization and invariants
One of the ways to determine if graphs are different is thorugh graph invariants. These are functions that evaluate to the same output if the graphs are the same under an equivalence relationship, i.e. isomorphism.
Definition A.18 (Graph invariant).
A graph invariant is a function $\xi$ on a graph that assigns the same value to isomorphic graphs. For a fixed $G$ and any other $G^{\prime}$ such that $G\simeq G^{\prime}$ we have $\xi(G)=\xi(G^{\prime})$. A graph invariant distinguishes non-isomorphic graphs $G$ and $G^{\prime}$, denoted $G\not\equiv_{\xi}G^{\prime}$ if $\xi(G)\neq\xi(G^{\prime})$ and conversely if $G$ and $G^{\prime}$ are isomorphic then $G\equiv_{\xi}G^{\prime}$ if $\xi(G)=\xi(G^{\prime})$.
An equivalent concept can be formulated in terms of combinatorial complexes, as we have a notion of equivalence from Definition˜A.17.
Definition A.19 (CC-Invariant).
A CC-Invariant is a function $\xi$ on a CC that assigns the same value to isomorphic CCs. For a fixed combinatorial complex $CC_{1}$ and any other $CC_{2}$ such that $CC_{1}\simeq CC_{2}$ we have $\xi(CC_{1})=\xi(CC_{2}^{\prime})$. As with graph invariants, this function helps distinguish non-isomorphic CCs.
Now notice that a CC-Invariant induced by a neighborhood tuple $(\mathcal{N}_{1},\mathcal{N}_{2})$ is simply a graph invariant on the resuling Hasse induced by the neighborhoods. With this in mind, Hasse graphs can be used as piece-wise invariants to compose an invariant over the whole complex.
Appendix B Theoretical motivation
Proposition B.1.
Let $\mathcal{CC}_{1}=(\mathcal{V}_{1},\mathcal{C}_{1},\mathrm{rk}_{1})$ and $\mathcal{CC}_{2}=(\mathcal{V}_{2},\mathcal{C}_{2},\mathrm{rk}_{2})$. Let $\mathcal{N}_{1}$ be a neighborhood function on $\mathcal{CC}_{1}$ and $\mathcal{N}_{2}$ be a neighborhood function on $\mathcal{CC}_{2}$. For every CC-Isomorphism $f:\mathcal{C}_{1}\to\mathcal{C}_{2}$ induced by $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, there exists a unique graph isomorphism between their respective strictly augmented Hasse graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
Proof.
Note that this is trivial since every isomorphism is also a homomorphism and by Proposition˜A.15 it is unique. ∎
First, we denote two special kinds of neighborhood functions. An incidence neighborhood function $\mathcal{N}$ is such that $\forall\sigma,\tau\in\mathcal{C}$, where $\sigma\in\mathcal{N}(\tau)$ or $\tau\in\mathcal{N}(\sigma)$ only one of these conditions is true:
  1. 1.
    $\mathrm{rk}_{1}(\sigma)>\mathrm{rk}_{1}(\tau)$,
  2. 2.
    $\mathrm{rk}_{1}(\sigma)<\mathrm{rk}_{1}(\tau)$.
Conversly, a adjacency neighborhood function $\mathcal{N}$ is such that $\forall\sigma,\tau\in\mathcal{C}$, where $\sigma\in\mathcal{N}(\tau)$ or $\tau\in\mathcal{N}(\sigma)$ then $\mathrm{rk}_{1}(\sigma)=\mathrm{rk}_{1}(\tau)$.
Proposition B.2.
Let $CC_{1}$ and $CC_{2}$ be two CCs where $CC_{1}\not\simeq CC_{2}$. There exists a pair of neighborhood functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, on $CC_{1}$ and $CC_{2}$ respectively, such that the induced Hasse graphs $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are non-isomorphic.
Proof.
Assume that for any choice of functions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ then $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are isomorphic. This means that there is an isomorphism $\hat{f}:\mathcal{G}_{\mathcal{N}_{1}}\to\mathcal{G}_{\mathcal{N}_{2}}$. This means that
$\forall u,v\in V(\mathcal{G}_{\mathcal{N}_{1}})\text{ s.t. }(u,v)\in E(\mathcal{G}_{\mathcal{N}_{1}})\implies(\hat{f}(u),\hat{f}(v))\in E(\mathcal{G}_{\mathcal{N}_{2}})$ (9)
and
$\forall u^{\prime},v^{\prime}\in V(\mathcal{G}_{\mathcal{N}_{2}})\text{ s.t. }(u^{\prime},v^{\prime})\in E(\mathcal{G}_{\mathcal{N}_{2}})\implies(\hat{f}^{-1}(u),\hat{f}^{-1}(v^{\prime}))\in E(\mathcal{G}_{\mathcal{N}_{1}}).$ (10)
Note that each isomorphism between Hasse graph also uniquely defines a CC-Isomorphism $f$ induced by the neighborhood functions $(\mathcal{N}_{1},\mathcal{N}_{2})$. Now, notice that only for a specific family of neighborhood functions, this maps encode rank information. This family are the incidence neighborhood functions of the form $\mathcal{I}_{s\to t}$ where $s\neq t$. Since $\mathcal{G}_{\mathcal{N}_{1}}$ and $\mathcal{G}_{\mathcal{N}_{2}}$ are directed graphs that means that rank is presserved in $f$. In other words, this implies that there is a CC-Isomorphism $g$ since for any restriction over the domain $\mathcal{C}_{1}$ induced by any arbitrary neighborhood function, there always exists a map to $\mathcal{C}_{2}$ where
$\forall\sigma,\tau\in\mathcal{C}_{1}\text{ if }\sigma\subseteq\tau\text{ then }f(\sigma)\subseteq f(\tau)$
holds trivially by selecting either an adjacency or an incidence since these neighborhood functions preserve set-composition and
$\forall\sigma\in\mathcal{C}_{1}$ (11)
holds since $f$ preserves the rank. However, by our initial construction $CC_{1}\not\simeq CC_{2}$ and thus we reach a contradiction. ∎
Appendix C Counting neighborhoods and message passing routes
(b) ${\mathcal{I}}_{1\to 0}$
(c) ${\mathcal{I}}_{2\to 0}$
(d) ${\mathcal{A}}_{0,1}$
(e) ${\mathcal{A}}_{0,2}$
(f) ${\mathcal{I}}_{0\to 1}$
(g) ${\mathcal{I}}_{2\to 1}$
(h) ${\mathcal{A}}_{1,0}$
(i) ${\mathcal{A}}_{1,2}$
(j) ${\mathcal{I}}_{0\to 2}$
(k) ${\mathcal{I}}_{1\to 2}$
(l) ${\mathcal{A}}_{2,0}$
(m) ${\mathcal{A}}_{2,1}$
Figure 3: Illustration of a combinatorial complex (top) and different neighborhood types, grouped by the target entity: nodes (second row), edges (third row), and faces (bottom row).
This section provides a detailed discussion of two key concepts in Topological Deep Learning: expanded neighborhoods and the combinatorial explosion of message passing routes. This discussion, while emphasizing the expressive power of combinatorial representations, highlights the challenges they pose when scaling to real-world problems.
C.1 Neighborhood types in combinatorial complexes
Combinatorial representations allow for a variety of neighborhood functions, depending on how cells of different ranks interact. For a 2-dimensional combinatorial complex (with 0-, 11-, and $2$-cells), Figure˜3 illustrates all possible neighborhood functions. These relations can be grouped into two main groups:
  • Incidence Relations:
    • (b) ${\mathcal{I}}_{1\to 0}$: edges $\rightarrow$ nodes
    • (c) ${\mathcal{I}}_{2\to 0}$: faces $\rightarrow$ nodes
    • (f) ${\mathcal{I}}_{0\to 1}$: nodes $\rightarrow$ edges
    • (g) ${\mathcal{I}}_{2\to 1}$: faces $\rightarrow$ edges
    • (j) ${\mathcal{I}}_{0\to 2}$: nodes $\rightarrow$ faces
    • (k) ${\mathcal{I}}_{1\to 2}$: edges $\rightarrow$ faces
  • Adjacency Relations:
    • (d) ${\mathcal{A}}_{0,1}$: nodes $\rightarrow$ nodes (via edges)
    • (e) ${\mathcal{A}}_{0,2}$: nodes $\rightarrow$ nodes (via faces)
    • (h) ${\mathcal{A}}_{1,0}$: edges $\rightarrow$ edges (via nodes)
    • (i) ${\mathcal{A}}_{1,2}$: edges $\rightarrow$ edges (via faces)
    • (l) ${\mathcal{A}}_{2,0}$: faces $\rightarrow$ faces (via nodes)
    • (m) ${\mathcal{A}}_{2,1}$: faces $\rightarrow$ faces (via edges)
C.2 Counting neighborhoods
The combinatorial complex depicted in Figure˜3 has rank $R=\dim({\mathcal{T}})=2$, which results in 12 distinct neighborhood functions. These neighborhood functions can be divided into the two categories from above: (i) incidence relations and (ii) adjacency relations.
Under these assumptions, the number of incidence relations is $|I|=(R+1)R$, as each of the $R+1$ ranks can be connected to any of the remaining $R$ ranks. Similarly, each rank gives rise to $R$ possible adjacency relations (each defined through a different intermediary rank), also resulting in $|A|=(R+1)R$. Combining these yields a total of
$|{\mathcal{N}}_{{\mathcal{T}}}|=2(R+1)R$
distinct neighborhood functions. For the example in Figure˜3, where $R=2$, this formula recovers the 12 neighborhoods observed.
C.3 Message passing routes
The number of neighborhood functions increases quadratically with the maximum rank, which may initially seem manageable, especially since most practical datasets involve low-rank structures (e.g., $R=2$ or $R=3$). However, the true combinatorial complexity arises when considering how the results neighborhoods can be composed. To formalize this notion, the following definition is introduced.
Definition C.1 (Message passing route).
A message passing route is an ordered set of neighborhood functions that a model uses to propagate information across a data structure.
Information may flow intra-rank, through adjacency relations, or inter-rank, through incidence relations. For example, a network might use the message passing route ${\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ to transmit information from faces to edges, then from edges to nodes, from nodes back to faces, and finally between nodes connected by an edge.
Given a fixed combinatorial complex, the number of possible message passing routes is infinite, since any sequence can reuse the same neighborhood relations and grow to arbitrary length. In what follows, the analysis is restricted to a subset of message passing routes that are particularly relevant for the design of higher-order neural architectures.
Definition C.2 (Minimal message passing route).
A minimal message passing route is a message passing route of minimal length that updates all ranks in a higher-order representation at least once. Furthermore, it ensures that, after a sufficient number of message passing steps, the representation of each rank depends on all other ranks.
Minimal routes are of particular interest, as they guarantee full interdependence between all cells—a desirable property for expressive models. Moreover, minimal length implies simplicity in design, yielding the most compact routes that preserve this full dependency. For instance, the previously introduced route ${\mathcal{I}}_{2\to 1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{0\to 2},{\mathcal{A}}_{0,1}]$ can be transformed into a minimal message passing route by removing the adjacency ${\mathcal{A}}_{0,1}$. In this reduced route, information from all cells becomes accessible to all others within two message passing steps. In fact, consider the case of 0-cells: after the first step, they depend on $1$-cells; after the second step, since $1$-cells now incorporate information from $2$-cells, the 0-cells become indirectly connected to them as well.
By definition, minimal message passing routes rely exclusively on incidence relations, as adjacencies do not propagate information inter-rank. Since each incidence targets a specific rank, and every rank must be updated once, minimal routes necessarily have length $R+1$. To enumerate all possible minimal routes, consider the set of ranks $T=\{0,1,\dots,R\}$, of size $R+1$. There are $(R+1)!$ distinct permutations of these ranks, each defining a unique order in which the ranks are updated. Given such a permutation $\{r_{0},r_{1},\dots,r_{R}\}$, a corresponding minimal message passing route can be constructed by selecting the sequence of incidence relations ${\mathcal{I}}_{r_{j-1}\to r_{j}}$ for $j=1,\dots,R$, and closing the cycle with ${\mathcal{I}}_{r_{R}\to r_{0}}$. This construction yields exactly one valid minimal route per permutation, resulting in a total of $N=(R+1)!$ such routes. In the case of graphs, where $R=1$, this results in exactly two minimal message passing routes.
C.4 Beyond minimal routes: including adjacencies
Minimal message passing routes represent only a small subset of the possible routes that can be constructed using the full set of neighborhoods in a combinatorial complex. A natural extension involves augmenting these minimal routes with adjacency operations, a common practice in many TDL architectures. For simplicity, the analysis here is restricted to the case where adjacencies are applied strictly after all incidence relations.
As previously discussed, for a given rank $r$, there are $R$ possible adjacency relations of the form ${\mathcal{A}}_{r,x}$. Since adjacencies operate intra-rank, they do not affect inter-rank dependencies and merely update the representation of the corresponding rank. As a consequence, the order in which these adjacencies are applied does not impact the outcome of the message passing route. For example, applying ${\mathcal{A}}_{0,0}$ followed by ${\mathcal{A}}_{1,1}$ produces the same result as applying them in reverse order.
Given that each of the $R+1$ ranks can be updated with any of their $R$ possible adjacencies, there are $(R+1)^{R}$ combinations of adjacency updates. When combined with the $(R+1)!$ minimal incidence-based routes, the total number of extended message passing routes becomes $\hat{N}=(R+1)!\cdot(R+1)^{R}$.
Appendix D Framework realization
This section contains an in-depth explanation of the empirical choices of PSEs that realize the theoretical results of this work, serving as a complement to Section˜4.2. It is important to recall that the previous theoretical results hold for any choice of functions $g_{k}(\cdot)$ and $\psi_{k}(\cdot)$ (characterization and aggregation, respectively), as long as they fulfill the injectivity and permutation equivariance constraints. The fact that a certain choice of functions does not completely fulfill the injectivity or permutation equivariance constraints does not imply that they have to be completely disregarded. These functions are still relevant to study the empirical reach of the theoretical results in a more constrained setting. More importantly, fully injective and permutation equivariant functions need to be learned (or fine-tuned) while training, greatly impacting the computational advantage of this method. As such, these endeavours are defered to future works.
D.1 Handcrafted PSEs
In the literature, functions over graphs that characterize their structure are commonly referred to as positional and structural encodings (PSEs). While it is true that not all PSEs are meant to act as fully characterizing functions, their purpose is to represent certain characteristics of the graph that are implicitly present in its structure. Constructing PSEs that fully characterize graphs up to isomorphism and preserve injectivity results in the graph canonization problem Dong et al. [2023]. Consequently, there will always be a trade-off between the goodness of the characterization and the computational complexity. With this in mind, the following choices of characterization functions stem from the literature as commonly used methods to augment the performance of GNNs. Additionally, while not fully injective, they are permutation equivariant, completely fulfilling that theoretical criterion.
Let ${\mathcal{G}}=(V,E)$ be an undirected graph with $|V|=n$ nodes. The characterization functions $g_{k}:{\mathcal{G}}\to\mathbb{R}^{d\times d}$ with arbitrary codomains used in this work are defined as follows:
  1. 1.
    $g_{1}(\cdot)=\mathrm{LapPE}_{i}=[\phi_{1},\dots,\phi_{i}]\in\mathbb{R}^{n\times i},$ the first $i$ nontrivial eigenvectors of the graph Laplacian (up to frequency $i$).
  2. 2.
    $g_{2}(\cdot)=\mathrm{RWSE}_{K}\in\mathbb{R}^{n\times K},$ the Random Walk Structural Encodings, obtained by stacking the diagonal entries of the tt‑step random‑walk transition matrices for $t=1,\dots,K$ (here $K=60$).
  3. 3.
    $g_{3}(\cdot)=\mathrm{HKdiagSE}_{K}\in\mathbb{R}^{n\times K},$ the Heat Kernel Diagonal Structural Encodings, given by $\exp(-\lambda_{j}\,t)]_{j=1,\dots,n;\,t=1,\dots,K}$ for the Laplacian eigenvalues $\{\lambda_{j}\}$.
  4. 4.
    $g_{4}(\cdot)=\mathrm{ElstaticPE}\in\mathbb{R}^{n\times d},$ the Electrostatic Potential Positional Encodings, computed by solving the discrete Poisson equation on $\mathcal{G}$} with suitable boundary conditions.
Importantly, since each $g_{k}(\cdot)$ depends only on the adjacency structure of $\mathcal{G}$}, none of these channels requires access to any node feature vectors. For more indepth explanation of each PSE, see Cantürk et al. [2023].
D.2 Pretrained encodings
As mentioned previously, this method relies on the use of GPSE. This is a pre-trained graph PSE presented by  Cantürk et al. [2023]. The pre-trained componenet consists of a GNN model pre-trained on a set of different datasets: ZINC, MolPCBA, GEOM, PCQM4Mv2 and ChEMBL. During the pre-training, a combination of losses targeting positional and structural properties is used to learn a latent representation of them.
In this case only one function $g_{1}=\text{GPSE}(\cdot)$ is used to calculate the encodings. It is worth noting that random noise initialization is used in GPSE to help the model learn some of the more challenging PEs. To emulate this procedure, for all neighborhoods ${\mathcal{N}}\in{\mathcal{N}}_{C}$, ${\mathcal{N}}_{C}\subseteq{\mathcal{N}}_{\mathcal{T}}$, and their corresponding Hasse graphs ${\mathcal{J}}_{{\mathcal{N}}}{\bigr|}_{k}=({\mathcal{C}}_{\mathcal{N}},{\mathcal{E}}_{\mathcal{N}},{\mathcal{F}}_{\mathcal{N}})$, the features $F_{r,k}\in{\mathcal{F}}_{\mathcal{N}}$ are initialized as $F_{r,k}\sim N(0,\mathbf{I})\in\mathbb{R}^{20}$.
D.3 Learnable embedding functions
For the functions $f_{r,k}$, $\varepsilon_{r}$, and $\Theta_{r}$ HOPSE uses $L$-layer multilayer perceptrons (MLPs) with independent parameters. Each layer calculates the next embedding as:
$\mathbf{h}_{r,k}^{(\ell)}=\sigma\!\Biggl(\mathrm{LayerNorm}\Bigl((\mathbf{h}_{r,k}^{(\ell-1)}\mathbf{W}_{r,k}^{(\ell)}+\mathbf{b}_{r,k}^{(\ell)})+\mathbf{h}_{r,k}^{(\ell-1)});\,\varepsilon=10^{-6}\Bigr)\Biggr),$ (12)
where each $\mathbf{W}_{r,k}^{(\ell)}\in\mathbb{R}^{D_{r}\times D_{r}}$ and $\mathbf{b}_{r,k}^{(\ell)}\in\mathbb{R}^{D_{r}}$ are learned, an additional $\mathbf{h}_{r,k}^{(\ell-1)}$ is added as a skip-connection, $\mathrm{LayerNorm}(\cdot;\varepsilon)$ normalizes to zero mean and unit variance with $\varepsilon=10^{-6}$ for numerical stability, and $\sigma$ is the pointwise nonlinearity LeakyReLU. The final output $\mathbf{h}_{r,k}^{(L)}$ yields the updated channel feature.
All MLPs share depth $L$ but may vary in hidden‐layer widths $\{d_{r,\ell}\}$. This unified design—with layer normalization before activation—ensures stable gradient flow and flexible expressivity across channels and ranks.
Appendix E Complexity analysis
The computational complexity of HOPSE scales linearly with the number of cells. This is possible by offloading the structure-dependent computations to a one-time preprocessing step. The resulting efficiency, confirmed by the experimental results in Table˜4, makes HOPSE well-suited for large-scale graph data. The following sections break down the process into two distinct parts (i.) preprocessing and (ii.) training time complexity.
Preprocessing overhead.
For positional and structural encodings, although the functions $g_{k}(\cdot)$ can, in principle, be any function (including trainable neural networks) the implementation is restricted to either graph-theoretic positional encodings or those generated by a pretrained GPSE model. This allows the tensors $\mathbf{X}_{r,k}$ (as defined in equation 2) to be pre-computed before training begins.
This is helpful because certain positional encodings can be computationally expensive. For instance, using the eigenvectors of the graph Laplacian requires eigendecomposition of a matrix $n\times n$, which has a complexity of ${\mathcal{O}}(n^{3})$. This makes such encodings impractical for large graphs if computed during training.
Train time computation.
The trainable components of the HOPSE framework are the functions $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$, for all $r=1,\dots,R$ and $k=1,\dots,K$. HOPSE M and G use MLPs with $L$ layers to model these functions.
These MLPs applied to $N$ elements with input and output dimensions $D_{\text{in}}$ and $D_{\text{out}}$ have a computational complexity of ${\mathcal{O}}(ND_{\text{in}}D_{\text{out}})$. For simplicity, let us consider fixed input and output dimensions for each rank. The complexity of generating the initial features $\hat{\mathbf{Z}}_{r}$ in Equation˜4 is
${\mathcal{O}}(N_{r}D_{r}^{2}L),$
where $N_{r}$ is the number of cells of rank $r$, and $D_{r}$ is the feature dimension.
Next, the cost of computing $\mathbf{X}_{r,k}$ for all $K$ positional encodings (as in Equation˜3) is:
${\mathcal{O}}(N_{r}D_{r}^{2}KL),$
assuming that each function $f_{r,k}$ preserves the feature dimensionality.
Assuming concatenation is used as the aggregation operation in Equation˜5, and that the output features have dimension $D_{r}$, the complexity of computing the final embeddings $\mathbf{H}_{r}$ is:
${\mathcal{O}}(N_{r}D_{r}^{2}(K+1)L).$
Combining all components, the total complexity for rank-$r$ cells is:
$C_{r}={\mathcal{O}}\left(N_{r}D_{r}^{2}L+N_{r}D_{r}^{2}KL+N_{r}D_{r}^{2}(K+1)L\right)={\mathcal{O}}(N_{r}D_{r}^{2}(K+1)L).$
The final complexity, assuming that all cells have the same input and hidden dimensions, is
$C_{\text{HOPSE}}=\sum_{r=1}^{R}C_{r}.$
This complexity scales linearly with the number of rank-$r$ cells and is independent of their connectivity. This is in sharp contrast with most higher-order neural architectures. For example, the Generalized Combinatorial Complex Neural Network (GCCN) analyzed in Papillon et al. [2024] has the following complexity:
$C_{\text{GCCN}}=\sum_{\mathcal{N}\in\mathcal{N}_{\mathcal{C}}}{\mathcal{O}}(\|\mathcal{N}\|_{0}D^{2}+\text{nrows}(\mathcal{N})d_{r^{\prime}}D+\text{nrows}(\mathcal{N}))+\sum_{r^{\prime}\in[0,R^{\prime}]}n_{r^{\prime}}n_{\mathcal{N}_{r^{\prime}}}D,$
where:
  • ${\mathcal{N}}_{C}$ is the set of neighborhoods considered,
  • $\mathcal{N}$ is a neighborhood from rank-$r$ to rank-$r^{\prime}$ cells,
  • $\|\mathcal{N}\|_{0}$ is the number of nonzero entries in $\mathcal{N}$,
  • $D$ is the feature dimension,
  • $d_{r^{\prime}}$ is the degree of rank-$r^{\prime}$ cells (assumed constant),
  • $n_{r^{\prime}}$ is the number of rank-$r^{\prime}$ cells,
  • $n_{\mathcal{N}_{r^{\prime}}}$ is the number of neighborhoods with target rank $r^{\prime}$.
Unlike HOPSE, the GCCN complexity depends on the number of cells and the structure of their relationships.
Appendix F Experimental configuration
All experiments were performed on a Linux system with 256 CPU cores, 1TB of RAM, and 8 NVIDIA A30 GPUs, each with 24GB of GPU memory. The implementation code and reproduction materials will be released upon manuscript acceptance.
F.1 Datasets
The datasets considered are five standard benchmarking datasets commonly used in the TDL literature, comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^6] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [Morris et al., 2020]. For regression, the ZINC dataset is used [Irwin et al., 2012]. Following standard practice in the TDL literature [Telyatnikov et al., 2024], liftings are applied to transform these graph datasets into higher-order domains. Specifically, the cycle-based lifting was used for the cell domain, and the clique complex lifting for the simplicial domain.
Additionally, the MANTRA Ballester et al. [2025] dataset comprises different triangulations of manifolds. A triangulation of a manifold $\mathcal{M}$} is a pair $(K,\iota)$ where $K$ is a simplicial complex and $\iota$ is a homeomorphism between $\mathcal{M}$} and the geometric realization of $K$. Concretely, triangulations encode higher-order topological information that the underlying graph, or 0-skeleton, does not encode. Since some triangulations are non-homeomorphic but share an underlying graph, the ability to distinguish triangulations requires that the model can access the topological information to make the correct predictions. This dataset emphasizes benchmarking a model’s predictive ability on purely topological properties. Thus, it aims to assess how much the model encodes topological information beyond the base graph. In contrast to the MANTRA paper, this work does not enhance node features with any additional manually derived features but instead initializes features for all cells as ones. No other datasets are included, such as the ones proposed in [Eitan et al., 2024], as they are subsumed by MANTRA. Since MANTRA contains a wider variety of homeomorphic features, categorizing them by homeomorphism class (NAME) or their Betti numbers is more general and complex.
Regarding $\beta_{0},\beta_{1},\beta_{2}$, following the Ballester et al. [2025] work, in the case of predicting Betti numbers, this work focuses on training a single model as a multivariate regression task to report corresponding results for each Betti number. The choice of using F1 over AUROC is based on class imbalances of the Betti numbers $\beta_{0},\beta_{1},\beta_{2}$ and of the predicted properties name (NAME) and orientability ORIENT. Additionally the task of $\beta_{0}$ is excluded as all graphs are connected thus rendering the label as $1$ for all cases.
F.2 Training and evaluation.
Each dataset splits into three parts, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively. The exception is ZINC, for which the predefined splits are used [Irwin et al., 2012]. The results are the average over 5 runs using different seeds.
Following the literature, different performance metrics are reported per dataset. Specifically, predictive accuracy is used for MUTAG, PROTEINS, NCI1, and NCI109. Mean absolute error (MAE) is used for ZINC. F1 score is used for MANTRA datasets. For each dataset, the mean and standard deviation of the chosen metric are computed across the five test sets.
Discrepancy in ZINC Results and other datasets
The discrepancies in our baseline results from common literature arise from our deliberate choice to use the TopoBench Telyatnikov et al. [2024] suite, a framework designed for standardized and fair comparison. Our goal was to evaluate all methods on an even playing field to rigorously assess the performance-efficiency trade-off. This standardization introduces several key differences compared to prior publications [Corso et al., 2020, Dwivedi et al., 2023, Jin et al., 2024, Bouritsas et al., 2022]:
  • Experimental protocol: A single 50/25/25 train/validation/test split and a constrained hyperparameter search are used for all models (please refer to in Section˜F.3). This differs from the k-fold cross-validation and extensive tuning found in other works.
  • Standardized components: TopoBench uses common encoders and readouts to fairly evaluate the core “backbone” of each model. This contrasts with studies using domain-specific knowledge, such as directly encoding molecular rings.
  • Feature handling: To ensure a fair comparison, TopoBench sometimes disregards certain features (e.g., edge attributes in ZINC or MUTAG) in favour of standardization, which impacts the final reported performance.
F.3 Hyperparameter search
The HOPSE architecture is optimized over eight variations in the neighborhood set. Each variation specifies a combination of neighborhoods. The considered neighborhood combinations are as follows:
  • ${\mathcal{A}}_{0,1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$, ${\mathcal{A}}_{2,1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{I}}_{0\to 1}$, ${\mathcal{I}}_{1\to 2}$
  • ${\mathcal{A}}_{0,1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{I}}_{0\to 1}$, ${\mathcal{I}}_{1\to 2}$, ${\mathcal{I}}_{1\to 0}$, ${\mathcal{I}}_{2\to 1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$, ${\mathcal{A}}_{1,0}$, ${\mathcal{A}}_{2,1}$, ${\mathcal{I}}_{0\to 1}$, ${\mathcal{I}}_{1\to 2}$, ${\mathcal{I}}_{1\to 0}$, ${\mathcal{I}}_{2\to 1}$
  • ${\mathcal{A}}_{0,1}$, ${\mathcal{A}}_{1,2}$, ${\mathcal{A}}_{0,2}$, ${\mathcal{A}}_{1,0}$, ${\mathcal{A}}_{2,1}$, ${\mathcal{A}}_{2,0}$
The first three configurations rely solely on adjacency-based neighborhoods. The following four integrate adjacency and incidence relations, while the final configuration includes jumps across ranks.
For both HOPSE-M and HOPSE-G, the set of neighborhood configurations described above is incorporated into the architecture search space, allowing the models to explore diverse neighborhoods.
HOPSE-G leverages the GPSE model to extract both positional and structural information. The evaluation is on four variants of the GPSE model, each pre-trained on a different dataset: ZINC, GEOM, MOLPCBA, and PCQM4MV2. These pre-trained models are treated as hyperparameters during the hyperparameter search. Overall, results show that GPSE models pre-trained on different datasets generalize well and consistently capture meaningful positional and structural features across all evaluated benchmarks, with no single pre-trained model consistently outperforming the others.
For the SaNN model, the number of hops needed to define both positional and structural encodings was treated as an additional hyperparameter.
Finally, the search space over architecture parameters is restricted to hyperparameters that are common across all models: SaNN, HOPSE, SCCNN, and SCN. A grid-search strategy is employed to identify the optimal parameters for each model-dataset combination. Specifically:
  • Encoder hidden dimension: {32, 64, 128, 256}
  • Model hidden dimensions: {64, 128, 256 }
  • Encoder dropout: {0, 0.25, 0.5}
  • Number of backbone layers: {1, 2, 3, 4}
  • Learning rate: {0.01, 0.001}
  • Weight decay: {0, 0.0001}
  • Batch size: {128, 256}
For SCN and SCCNN models, the considered readout types are direct readout (DR) and signal down-propagation (SDP) (refer to Telyatnikov et al. [2024] for in-depth discussion on these options).
Appendix G Complete model performance and runtimes
Table 3: Performance on benchmark datasets.
Graph Tasks Topological Tasks Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC-12K ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 69.79 $\pm$ 6.80 75.70 $\pm$ 2.14 72.86 $\pm$ 0.69 72.20 $\pm$ 1.22 0.62 $\pm$ 0.01 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 72.77 $\pm$ 2.77 76.34 $\pm$ 1.66 75.00 $\pm$ 0.99 73.80 $\pm$ 0.73 0.61 $\pm$ 0.01 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 79.57 $\pm$ 6.13 75.20 $\pm$ 3.30 74.26 $\pm$ 0.96 74.42 $\pm$ 0.70 0.57 $\pm$ 0.04 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 73.62 $\pm$ 6.13 75.27 $\pm$ 2.14 74.49 $\pm$ 1.03 75.70 $\pm$ 1.04 0.53 $\pm$ 0.04 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 76.17 $\pm$ 6.63 74.19 $\pm$ 2.86 76.60 $\pm$ 1.75 77.12 $\pm$ 1.07 0.36 $\pm$ 0.02 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 79.57 $\pm$ 3.56 75.05 $\pm$ 3.52 68.75 $\pm$ 1.64 68.17 $\pm$ 0.88 0.55 $\pm$ 0.01 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 85.96 $\pm$ 4.66 74.91 $\pm$ 2.51 76.85 $\pm$ 1.50 75.76 $\pm$ 1.28 0.35 $\pm$ 0.01 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M (Our) 86.38 $\pm$ 6.31 75.13 $\pm$ 3.66 77.35 $\pm$ 0.81 76.34 $\pm$ 1.40 0.39 $\pm$ 0.02 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G (Our) 85.53 $\pm$ 2.77 76.27 $\pm$ 2.10 77.70 $\pm$ 1.40 76.36 $\pm$ 1.75 0.43 $\pm$ 0.01 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25 Cell CCCN 77.02 $\pm$ 9.32 73.33 $\pm$ 2.30 76.67 $\pm$ 1.48 75.35 $\pm$ 1.50 0.34 $\pm$ 0.02 - - - - CWN 80.43 $\pm$ 1.78 76.13 $\pm$ 2.70 73.93 $\pm$ 1.87 73.80 $\pm$ 2.06 0.34 $\pm$ 0.01 - - - - GCCN 86.38 $\pm$ 6.49 74.41 $\pm$ 1.77 78.23 $\pm$ 1.47 77.19 $\pm$ 0.21 0.19 $\pm$ 0.00 - - - - HOPSE-M (Our) 88.09 $\pm$ 5.55 76.99 $\pm$ 2.63 78.09 $\pm$ 1.44 77.27 $\pm$ 1.01 0.24 $\pm$ 0.01 - - - - HOPSE-G (Our) 88.09 $\pm$ 1.90 76.42 $\pm$ 2.19 79.07 $\pm$ 0.65 77.81 $\pm$ 1.26 0.21 $\pm$ 0.01 - - - -
Table 4: Total end-to-end training runtime in seconds.
Graph Tasks Topological Tasks Model MUTAG PROTEINS NCI1 NCI109 ZINC-12K NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ Graph GCN 4 $\pm$ 1 8 $\pm$ 2 53 $\pm$ 20 37 $\pm$ 9 147 $\pm$ 28 3538 $\pm$ 38 3534 $\pm$ 49 7250 $\pm$ 234 GAT 4 $\pm$ 1 8 $\pm$ 2 57 $\pm$ 17 56 $\pm$ 9 171 $\pm$ 65 3589 $\pm$ 84 3544 $\pm$ 44 3612 $\pm$ 52 GIN 5 $\pm$ 1 9 $\pm$ 2 61 $\pm$ 24 50 $\pm$ 8 168 $\pm$ 107 3568 $\pm$ 44 3536 $\pm$ 46 3674 $\pm$ 176 Simplicial SCN 8 $\pm$ 2 43 $\pm$ 13 276 $\pm$ 63 226 $\pm$ 66 1210 $\pm$ 572 5115 $\pm$ 653 4773 $\pm$ 193 6992 $\pm$ 2049 SCCNN 14 $\pm$ 3 54 $\pm$ 11 307 $\pm$ 83 354 $\pm$ 106 2060 $\pm$ 408 5521 $\pm$ 319 6920 $\pm$ 1524 6720 $\pm$ 210 SaNN 9 $\pm$ 2 50 $\pm$ 12 250 $\pm$ 18 211 $\pm$ 66 1263 $\pm$ 252 5572 $\pm$ 1053 4619 $\pm$ 103 5220 $\pm$ 421 GCCN 61 $\pm$ 7 66 $\pm$ 21 904 $\pm$ 180 538 $\pm$ 39 3603 $\pm$ 475 5898 $\pm$ 1961 4564 $\pm$ 631 7689 $\pm$ 2522 HOPSE-M (Ours) 7 $\pm$ 2 38 $\pm$ 8 164 $\pm$ 79 219 $\pm$ 26 1061 $\pm$ 405 3596 $\pm$ 519 3409 $\pm$ 162 5377 $\pm$ 185 HOPSE-G (Ours) 5 $\pm$ 0 33 $\pm$ 8 130 $\pm$ 22 107 $\pm$ 28 478 $\pm$ 107 3324 $\pm$ 1113 2795 $\pm$ 112 3876 $\pm$ 1141 Cell CCCN 12 $\pm$ 2 42 $\pm$ 7 372 $\pm$ 109 272 $\pm$ 21 1621 $\pm$ 142 - - - CWN 11 $\pm$ 1 54 $\pm$ 18 302 $\pm$ 63 295 $\pm$ 46 1390 $\pm$ 97 - - - GCCN 61 $\pm$ 18 59 $\pm$ 18 523 $\pm$ 119 386 $\pm$ 76 3301 $\pm$ 440 - - - HOPSE-M (Ours) 7 $\pm$ 1 37 $\pm$ 7 147 $\pm$ 24 122 $\pm$ 18 1313 $\pm$ 132 - - - HOPSE-G (Ours) 7 $\pm$ 1 32 $\pm$ 5 153 $\pm$ 8 104 $\pm$ 38 1230 $\pm$ 527 - - -
Appendix H Performance ablation across different neighborhoods
The choice of neighborhood functions set ${\mathcal{N}}_{C}$ is a crucial and non-trivial choice for the performance of HOPSE. This section contains the performance of HOPSE for each choice of neighborhood encoded by a distinct name, as reported in Table˜5.
Table 5: Taxonomy of neighborhood combinations
Group Neighborhood Combination
Adj-1 ${\mathcal{A}}_{0,1}$
Adj-2 ${\mathcal{A}}_{0,1},A_{1,2}$
Adj-3 ${\mathcal{A}}_{0,1},A_{1,2},A_{2,1}$
Inc-1 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2}$
Inc-2 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
Inc-3 ${\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
Mix-1 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{1\to 2},{\mathcal{I}}_{1\to 0},{\mathcal{I}}_{2\to 1}$
Mix-2 ${\mathcal{A}}_{0,1},{\mathcal{A}}_{1,2},{\mathcal{A}}_{0,2},{\mathcal{A}}_{1,0},{\mathcal{A}}_{2,1},{\mathcal{A}}_{2,0}$
For small and simple datasets (MUTAG, PROTEINS, NCI1, NCI109), the performance differences between neighborhood configurations are minimal, with most results falling within one standard deviation of each other. As datasets grow bigger and more complex, such as ZINC-12K and the MANTRA, the neighborhood choice starts to matter. The results show that for simplicial models, the best performing neighborhoods for ZINC-12K are the extended groups of adjacencies Adj-2 and Adj-3. However, for the MANTRA tasks, the most important neighborhood is Mix-2.Note that the per neighborhood ablation results differ from the main table due to some neighborhoods having a higher validation which does not lead to ahigher test.These findings demonstrate that the choice of neighborhood configuration represents a critical design decision in the topological learning frameworks. While simpler configurations may suffice for some graph benchmarks, an effective strategy is to consider a greater neighbourhood set ${\mathcal{N}}_{C}$ to capture the full expressive power of topological structures.
Table 6: Performance ablation of $\mathcal{N}_{S}$ on simplicial domain.
Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) HOPSE-M Adj-1 $83.40\pm 0.07$.12) $76.13\pm 0.03$.09) $75.97\pm 0.01$.18) $75.70\pm 0.01$.17) $0.43\pm 0.00$.01) $85.36\pm 0.08$.46) $83.76\pm 0.00$.94) $90.32\pm 0.00$.09) $68.79\pm 0.00$.78) Adj-2 $85.96\pm 0.05$.55) $76.85\pm 0.01$.75) $76.48\pm 0.01$.44) $75.37\pm 0.00$.98) $0.42\pm 0.00$.03) $94.45\pm 0.00$.82) $89.24\pm 0.00$.22) $91.72\pm 0.01$.17) $78.45\pm 0.01$.92) Adj-3 $84.68\pm 0.05$.51) $76.13\pm 0.03$.01) $75.64\pm 0.00$.76) $75.06\pm 0.00$.84) $0.43\pm 0.00$.01) $96.01\pm 0.00$.54) $88.60\pm 0.00$.77) $91.13\pm 0.01$.20) $79.12\pm 0.01$.11) Inc-1 $84.26\pm 0.05$.55) $76.06\pm 0.03$.05) $75.51\pm 0.01$.03) $75.20\pm 0.00$.93) $0.42\pm 0.00$.01) $96.27\pm 0.00$.28) $87.34\pm 0.00$.74) $90.76\pm 0.01$.57) $78.32\pm 0.01$.13) Inc-2 $85.11\pm 0.07$.22) $75.27\pm 0.01$.88) $75.51\pm 0.00$.53) $74.42\pm 0.00$.80) $0.43\pm 0.00$.02) $95.96\pm 0.00$.29) $86.35\pm 0.02$.15) $91.22\pm 0.00$.06) $79.08\pm 0.00$.75) Inc-3 $84.68\pm 0.06$.96) $76.13\pm 0.01$.84) $76.01\pm 0.00$.64) $74.77\pm 0.01$.21) $0.42\pm 0.00$.01) $96.07\pm 0.00$.19) $85.74\pm 0.01$.03) $91.07\pm 0.00$.10) $77.73\pm 0.01$.69) Mix-1 $82.98\pm 0.03$.98) $76.56\pm 0.02$.84) $75.66\pm 0.00$.32) $75.84\pm 0.00$.50) $0.43\pm 0.00$.01) $93.84\pm 0.01$.76) $87.42\pm 0.00$.64) $91.63\pm 0.00$.93) $80.45\pm 0.01$.06) Mix-2 $83.83\pm 0.04$.90) $76.06\pm 0.02$.31) $75.21\pm 0.01$.30) $75.90\pm 0.00$.48) $0.44\pm 0.00$.01) $95.71\pm 0.01$.41) $89.23\pm 0.00$.15) $90.87\pm 0.01$.07) $82.29\pm 0.00$.46) HOPSE-G Adj-1 $82.98\pm 0.07$.22) $74.05\pm 0.02$.17) $77.18\pm 0.01$.45) $76.21\pm 0.00$.65) $0.44\pm 0.00$.01) $80.43\pm 0.01$.03) $60.92\pm 0.01$.18) $88.18\pm 0.00$.03) $24.59\pm 0.01$.87) Adj-2 $85.53\pm 0.02$.77) $74.05\pm 0.03$.08) $77.51\pm 0.01$.15) $74.91\pm 0.01$.17) $43\pm 0.00.00$) $80.19\pm 0.01$.13) $61.88\pm 0.00$.83) $88.13\pm 0.00$.02) $20.31\pm 0.02$.57) Adj-3 $85.11\pm 0.03$.36) $73.26\pm 0.04$.20) $77.32\pm 0.01$.19) $74.79\pm 0.00$.91) $0.43\pm 0.00$.01) $79.11\pm 0.01$.53) $61.86\pm 0.01$.10) $88.11\pm 0.00$.02) $19.08\pm 0.03$.57) Inc-1 $84.68\pm 0.03$.50) $73.19\pm 0.03$.52) $77.70\pm 0.01$.40) $76.36\pm 0.01$.75) $44\pm 0.00.00$) $80.70\pm 0.00$.80) $62.17\pm 0.00$.98) $87.31\pm 0.01$.17) $0.47\pm 0.03.22$) Inc-2 $82.13\pm 0.05$.75) $76.27\pm 0.02$.10) $75.66\pm 0.01$.52) $74.17\pm 0.01$.99) $0.45\pm 0.00$.02) $78.92\pm 0.01$.25) $59.57\pm 0.01$.56) $88.08\pm 0.00$.08) $20.28\pm 0.03$.48) Inc-3 $77.87\pm 0.04$.90) $75.48\pm 0.03$.15) $6.15\pm 0.01.77$) $75.14\pm 0.00$.75) $0.44\pm 0.00$.01) $78.78\pm 0.00$.74) $60.35\pm 0.01$.05) $88.12\pm 0.00$.04) $18.82\pm 0.03$.54) Mix-1 $79.15\pm 0.07$.88) $75.70\pm 0.03$.24) $76.50\pm 0.01$.09) $75.35\pm 0.00$.97) $0.45\pm 0.00$.01) $79.76\pm 0.00$.51) $60.33\pm 0.01$.60) $88.07\pm 0.00$.09) $23.48\pm 0.03$.44) Mix-2 $83.40\pm 0.08$.56) $74.05\pm 0.03$.39) $77.02\pm 0.00$.92) $74.75\pm 0.01$.47) $0.46\pm 0.00$.01) $81.75\pm 0.01$.26) $61.32\pm 0.02$.67) $88.28\pm 0.00$.08) $35.37\pm 0.02$.25)
Table 7: Performance ablation of $\mathcal{N}_{S}$ on cell domain.
Graph Tasks $\mathcal{N}_{S}$ MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) HOPSE-M Adj-1 $83.40\pm 0.04$.85) $75.05\pm 0.02$.33) $76.09\pm 0.01$.99) $76.38\pm 0.00$.73) $0.27\pm 0.00$.01) Adj-2 $83.83\pm 0.03$.87) $77.28\pm 0.02$.44) $76.71\pm 0.00$.42) $75.80\pm 0.01$.71) $0.26\pm 0.00$.01) Adj-3 $85.53\pm 0.04$.85) $76.06\pm 0.03$.24) $76.05\pm 0.01$.75) $75.68\pm 0.00$.93) $25\pm 0.00.00$) Inc-1 $83.83\pm 0.05$.12) $76.99\pm 0.01$.94) $77.88\pm 0.01$.71) $76.32\pm 0.01$.40) $0.26\pm 0.00$.01) Inc-2 $86.38\pm 0.05$.34) $75.84\pm 0.02$.94) $76.63\pm 0.01$.78) $76.03\pm 0.01$.45) $0.26\pm 0.00$.01) Inc-3 $80.43\pm 0.05$.51) $75.70\pm 0.03$.55) $76.77\pm 0.00$.21) $75.78\pm 0.01$.66) $0.25\pm 0.00$.02) Mix-1 $84.26\pm 0.06$.49) $76.63\pm 0.01$.99) $77.80\pm 0.00$.82) $77.08\pm 0.01$.35) $0.24\pm 0.00$.01) Mix-2 $83.83\pm 0.04$.15) $77.06\pm 0.02$.43) $76.52\pm 0.02$.26) $76.22\pm 0.01$.64) $0.24\pm 0.00$.01) HOPSE-G Adj-1 $82.13\pm 0.07$.76) $76.27\pm 0.02$.70) $78.31\pm 0.00$.84) $77.48\pm 0.00$.99) $26\pm 0.00.00$) Adj-2 $87.66\pm 0.06$.96) $76.42\pm 0.02$.19) $78.83\pm 0.01$.40) $77.48\pm 0.00$.64) $24\pm 0.00.00$) Adj-3 $88.09\pm 0.01$.90) $75.91\pm 0.01$.89) $78.35\pm 0.01$.11) $76.94\pm 0.01$.46) $23\pm 0.00.00$) Inc-1 $85.11\pm 0.06$.56) $74.62\pm 0.03$.99) $77.57\pm 0.01$.67) $76.81\pm 0.01$.23) $0.25\pm 0.00$.01) Inc-2 $82.55\pm 0.08$.82) $74.91\pm 0.02$.11) $78.15\pm 0.00$.75) $76.40\pm 0.01$.86) $0.27\pm 0.00$.01) Inc-3 $4.68\pm 0.07.88$) $74.12\pm 0.03$.39) $76.81\pm 0.00$.59) $75.26\pm 0.01$.16) $26\pm 0.00.00$) Mix-1 $88.09\pm 0.03$.23) $76.20\pm 0.02$.62) $79.03\pm 0.00$.73) $76.88\pm 0.00$.95) $23\pm 0.00.00$) Mix-2 $85.11\pm 0.06$.02) $75.63\pm 0.02$.35) $79.07\pm 0.00$.65) $77.81\pm 0.01$.26) $0.21\pm 0.00$.01)
Appendix I Runtime ablation across different neighborhoods
I.1 Total runtime
On HOPSE-G, some more complex neighborhoods will result in faster time, even though they contain more information and thus parameters to be learned. This is one of the effects of convergence speed. This can be seen in neighborhoods Mix-1 and Inc-[1,3]. On HOPSE-M, there is no such drastic difference in HOPSE configurations, especially with mixed neighborhood sets, which provide an excellent balance between efficiency and expressiveness.
Table 8: Total training runtime (seconds) ablation of $\mathcal{N}_{S}$ on simplicial domain.
Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 $9.62\pm 0.00$.40) $49.15\pm 0.39$.24) $160.18\pm 0.29$.76) $145.63\pm 0.13$.40) $1488.50\pm 2.83$.52) $1754.92\pm 11.74$.29) $2281.97\pm 4.78$.17) $2498.81\pm 1.86$.09) Adj-2 $10.97\pm 0.01$.91) $49.32\pm 0.16$.14) $141.54\pm 0.34$.29) $177.84\pm 0.49$.45) $2855.15\pm 4.79$.89) $2455.75\pm 4.27$.45) $2478.94\pm 5.20$.52) $2331.82\pm 3.46$.01) Adj-3 $12.66\pm 0.02$.91) $47.95\pm 0.23$.11) $225.36\pm 0.45$.43) $195.86\pm 0.36$.31) $599.04\pm 0.42.11$) $3576.27\pm 3.02$.11) $3043.11\pm 8.31$.36) $2561.17\pm 1.63$.73) Inc-1 $11.50\pm 0.02$.44) $58.85\pm 0.13$.74) $169.46\pm 0.42$.22) $220.22\pm 0.32$.95) $2655.44\pm 4.36$.38) $2963.63\pm 5.10$.20) $2287.67\pm 3.48$.83) $2613.46\pm 1.72$.92) Inc-2 $10.65\pm 0.01$.13) $44.15\pm 0.20$.68) $162.13\pm 0.44$.12) $139.93\pm 0.22$.47) $1923.95\pm 3.35$.87) $3287.92\pm 4.24$.70) $2439.12\pm 4.94$.99) $2178.14\pm 3.33$.52) Inc-3 $11.20\pm 0.01$.06) $61.33\pm 0.21$.24) $236.34\pm 0.21$.93) $194.27\pm 0.37$.37) $2620.02\pm 7.80$.01) $3292.73\pm 4.44$.17) $2825.51\pm 4.50$.02) $2386.66\pm 3.17$.80) Mix-1 $12.16\pm 0.00$.93) $74.15\pm 0.14$.39) $372.57\pm 0.43$.27) $235.41\pm 0.44$.74) $3168.39\pm 5.57$.44) $4205.60\pm 15.93$.84) $6681.04\pm 12.61$.16) $3591.07\pm 4.58$.88) Mix-2 $11.17\pm 0.00$.84) $57.38\pm 0.19$.99) $210.32\pm 0.53$.70) $217.65\pm 0.27$.40) $1630.30\pm 0.75$.76) $3497.96\pm 11.59$.55) $3737.92\pm 2.34$.34) $3590.55\pm 4.68$.47) HOPSE-G Adj-1 $4.56\pm 0.00$.62) $25.49\pm 0.06$.16) $92.97\pm 0.27$.48) $104.94\pm 0.28$.21) $542.03\pm 0.94$.68) $1891.07\pm 0.17$.81) $1885.35\pm 0.28$.06) $2194.20\pm 0.42$.70) Adj-2 $4.92\pm 0.00$.37) $32.03\pm 0.09$.91) $118.32\pm 0.17$.84) $111.18\pm 0.18$.27) $480.02\pm 1.00$.70) $2172.54\pm 2.16$.62) $2096.15\pm 0.33$.94) $2323.75\pm 2.22$.16) Adj-3 $5.58\pm 0.01$.28) $39.85\pm 0.17$.90) $120.88\pm 0.31$.47) $134.65\pm 0.49$.81) $478.06\pm 1.07$.13) $2302.00\pm 0.47$.08) $2385.63\pm 1.76$.04) $2997.61\pm 8.24$.53) Inc-1 $5.80\pm 0.00$.27) $32.71\pm 0.08$.43) $130.06\pm 0.22$.37) $106.98\pm 0.27$.98) $649.05\pm 1.14$.28) $2580.45\pm 0.47$.63) $2471.89\pm 0.47$.46) $3716.82\pm 11.92$.83) Inc-2 $6.44\pm 0.00$.54) $32.98\pm 0.08$.25) $128.01\pm 0.22$.39) $99.39\pm 0.33$.07) $661.57\pm 1.33$.52) $2489.63\pm 4.44$.47) $2240.50\pm 0.43$.84) $2568.75\pm 0.76$.80) Inc-3 $6.85\pm 0.00$.98) $51.18\pm 0.11$.23) $174.20\pm 0.57$.02) $148.67\pm 0.37$.09) $778.78\pm 1.17$.23) $3038.25\pm 0.65$.40) $2930.43\pm 4.30$.59) $3602.56\pm 5.55$.32) Mix-1 $6.85\pm 0.00$.72) $51.94\pm 0.09$.59) $209.07\pm 0.28$.40) $150.07\pm 0.36$.64) $911.69\pm 1.65$.22) $3408.18\pm 5.15$.16) $3379.79\pm 1.11$.99) $4321.64\pm 10.94$.92) Mix-2 $5.16\pm 0.00$.36) $35.97\pm 0.04$.38) $168.30\pm 0.38$.75) $142.36\pm 0.34$.54) $716.46\pm 0.53$.19) $3324.16\pm 11.13$.48) $2794.73\pm 3.45$.73) $3876.34\pm 11.41$.15)
Table 9: Total training runtime (seconds) of $\mathcal{N}_{S}$ on cell domain.
Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K HOPSE-M Adj-1 $8.61\pm 0.02$.06) $38.36\pm 0.12$.45) $153.99\pm 0.20$.76) $155.94\pm 0.40$.02) $856.03\pm 2.24$.92) Adj-2 $5.51\pm 0.00$.81) $32.18\pm 0.10$.10) $169.57\pm 0.35$.97) $153.25\pm 0.21$.41) $1013.70\pm 1.20$.74) Adj-3 $5.50\pm 0.00$.41) $33.48\pm 0.06$.54) $175.46\pm 0.22$.23) $153.43\pm 0.17$.47) $922.75\pm 0.89$.20) Inc-1 $6.98\pm 0.01$.24) $41.18\pm 0.15$.30) $144.62\pm 0.29$.66) $192.74\pm 0.45$.31) $952.75\pm 1.66$.32) Inc-2 $6.81\pm 0.00$.88) $36.17\pm 0.09$.11) $154.09\pm 0.55$.24) $135.15\pm 0.29$.51) $827.17\pm 0.77$.46) Inc-3 $7.15\pm 0.01$.07) $44.15\pm 0.16$.85) $234.99\pm 0.25$.73) $157.33\pm 0.30$.17) $934.61\pm 0.49$.49) Mix-1 $8.88\pm 0.01$.71) $30.30\pm 0.02$.29) $163.48\pm 0.27$.28) $222.96\pm 0.59$.41) $1421.85\pm 1.27$.42) Mix-2 $5.54\pm 0.00$.54) $36.30\pm 0.08$.16) $129.82\pm 0.21$.18) $161.53\pm 0.50$.71) $1046.86\pm 0.47$.16) HOPSE-G Adj-1 $5.09\pm 0.00$.93) $26.71\pm 0.04$.79) $88.31\pm 0.11$.27) $95.95\pm 0.28$.16) $683.86\pm 1.25$.70) Adj-2 $4.98\pm 0.00$.47) $31.85\pm 0.05$.43) $96.33\pm 0.19$.91) $93.98\pm 0.15$.00) $887.83\pm 1.60$.74) Adj-3 $6.64\pm 0.00$.89) $31.13\pm 0.05$.56) $100.65\pm 0.14$.14) $113.39\pm 0.18$.75) $595.39\pm 1.07$.46) Inc-1 $5.75\pm 0.01$.45) $34.11\pm 0.07$.89) $110.19\pm 0.19$.33) $105.67\pm 0.27$.00) $767.42\pm 0.85$.39) Inc-2 $7.65\pm 0.02$.32) $24.65\pm 0.06$.74) $105.02\pm 0.28$.61) $126.14\pm 0.37$.05) $721.57\pm 0.60$.10) Inc-3 $7.87\pm 0.01$.47) $44.42\pm 0.09$.81) $181.18\pm 0.96$.45) $100.22\pm 0.11$.50) $896.86\pm 1.62$.01) Mix-1 $7.61\pm 0.00$.68) $46.14\pm 0.10$.02) $239.59\pm 0.38$.24) $168.35\pm 0.50$.01) $671.40\pm 9.48$.29) Mix-2 $5.27\pm 0.00$.57) $36.29\pm 0.13$.05) $152.73\pm 0.08$.26) $104.12\pm 0.37$.76) $1229.56\pm 5.27$.27)
I.2 Runtime per epoch
Tables˜10 and 11 validate that smaller neighborhoods, and thus less parameters, result in better per epoch runtimes. HOPSE-G has lower per epoch runtime due to having to learn less instantiations of $f_{r,k}$ since $k=1$. Simpler configurations such as Adj-1 consistently yield the lowest runtimes, particularly for HOPSE-M, where Adj-1 achieves the fastest epoch times in 5 out of 8 datasets. Note that as dataset size increases, the epoch time scales linearly, thus while HOPSE-G remains the fastest, it is only so by a increasingly smaller fraction.
Table 10: Per epoch runtime (seconds) of $\mathcal{N}_{S}$ on simplicial domain.
Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 $03\pm 0.00.00$) $0.12\pm 0.00$.04) $0.40\pm 0.00$.01) $0.34\pm 0.00$.01) $0.43\pm 0.00$.03) $3.12\pm 0.00$.11) $3.67\pm 0.00$.07) $3.91\pm 0.00$.14) Adj-2 $03\pm 0.00.00$) $0.12\pm 0.00$.01) $0.44\pm 0.00$.01) $0.37\pm 0.00$.02) $0.89\pm 0.00$.07) $3.52\pm 0.00$.10) $4.15\pm 0.00$.09) $4.20\pm 0.00$.24) Adj-3 $03\pm 0.00.00$) $0.15\pm 0.00$.07) $0.54\pm 0.00$.01) $0.39\pm 0.00$.02) $0.62\pm 0.00$.11) $4.67\pm 0.00$.06) $4.40\pm 0.00$.12) $4.36\pm 0.00$.12) Inc-1 $03\pm 0.00.00$) $0.15\pm 0.00$.01) $0.51\pm 0.00$.02) $0.47\pm 0.00$.06) $0.88\pm 0.00$.06) $5.20\pm 0.00$.24) $4.33\pm 0.00$.12) $4.83\pm 0.00$.12) Inc-2 $03\pm 0.00.00$) $0.14\pm 0.00$.02) $0.45\pm 0.00$.01) $0.44\pm 0.00$.02) $0.64\pm 0.00$.05) $4.58\pm 0.00$.07) $4.18\pm 0.00$.10) $4.39\pm 0.00$.19) Inc-3 $03\pm 0.00.00$) $0.17\pm 0.00$.01) $0.54\pm 0.00$.02) $0.48\pm 0.00$.04) $0.87\pm 0.00$.06) $5.62\pm 0.00$.12) $5.43\pm 0.00$.09) $5.80\pm 0.00$.18) Mix-1 $04\pm 0.00.00$) $0.22\pm 0.00$.02) $0.55\pm 0.00$.01) $0.55\pm 0.00$.01) $1.23\pm 0.00$.13) $5.94\pm 0.00$.32) $8.08\pm 0.00$.15) $6.64\pm 0.00$.32) Mix-2 $03\pm 0.00.00$) $0.15\pm 0.00$.01) $0.50\pm 0.00$.02) $0.46\pm 0.00$.01) $0.61\pm 0.00$.04) $4.82\pm 0.00$.24) $4.72\pm 0.00$.26) $5.65\pm 0.00$.42) HOPSE-G Adj-1 0.02(0.00) $0.10\pm 0.00$.01) 0.28(0.01) $28\pm 0.00.00$) $0.26\pm 0.00$.02) $3.19\pm 0.00$.09) $3.18\pm 0.00$.12) $3.55\pm 0.00$.12) Adj-2 $02\pm 0.00.00$) $0.12\pm 0.00$.01) $30\pm 0.00.00$) $0.31\pm 0.00$.01) $0.28\pm 0.00$.01) $3.53\pm 0.00$.12) $3.41\pm 0.00$.10) $3.67\pm 0.00$.12) Adj-3 $02\pm 0.00.00$) $11\pm 0.00.00$) $0.33\pm 0.00$.02) $0.34\pm 0.00$.01) $0.29\pm 0.00$.02) $3.79\pm 0.00$.14) $3.84\pm 0.00$.14) $3.98\pm 0.00$.07) Inc-1 $02\pm 0.00.00$) $0.13\pm 0.00$.01) $0.38\pm 0.00$.01) $0.37\pm 0.00$.02) $0.34\pm 0.00$.01) $4.44\pm 0.00$.14) $4.23\pm 0.00$.12) $4.64\pm 0.00$.08) Inc-2 $02\pm 0.00.00$) $10\pm 0.00.00$) $31\pm 0.00.00$) $31\pm 0.00.00$) $0.28\pm 0.00$.01) $3.74\pm 0.00$.17) $3.66\pm 0.00$.10) $3.78\pm 0.00$.06) Inc-3 $02\pm 0.00.00$) $13\pm 0.00.00$) $41\pm 0.00.00$) $0.40\pm 0.00$.01) $0.38\pm 0.00$.03) $5.17\pm 0.00$.40) $4.75\pm 0.00$.23) $5.13\pm 0.00$.06) Mix-1 $02\pm 0.00.00$) $0.16\pm 0.00$.01) $0.46\pm 0.00$.02) $0.46\pm 0.00$.01) $0.44\pm 0.00$.01) $5.50\pm 0.00$.24) $5.74\pm 0.00$.22) $5.84\pm 0.00$.38) Mix-2 $02\pm 0.00.00$) $0.13\pm 0.00$.01) $38\pm 0.00.00$) $0.38\pm 0.00$.01) $0.36\pm 0.00$.02) $4.48\pm 0.00$.14) $4.55\pm 0.00$.19) $4.84\pm 0.00$.11)
Table 11: Per epoch runtime (seconds) of $\mathcal{N}_{S}$ on cell domain.
Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC-12K HOPSE-M Adj-1 $02\pm 0.00.00$) $12\pm 0.00.00$) $0.39\pm 0.00$.01) $0.44\pm 0.00$.05) $0.27\pm 0.00$.01) Adj-2 $02\pm 0.00.00$) $0.10\pm 0.00$.01) $0.29\pm 0.00$.01) $0.38\pm 0.00$.02) $0.27\pm 0.00$.01) Adj-3 $02\pm 0.00.00$) $0.11\pm 0.00$.01) $0.36\pm 0.00$.01) $0.42\pm 0.00$.02) $0.30\pm 0.00$.01) Inc-1 $02\pm 0.00.00$) $0.12\pm 0.00$.01) $0.40\pm 0.00$.01) $0.47\pm 0.00$.02) $0.44\pm 0.00$.01) Inc-2 $03\pm 0.00.00$) $0.12\pm 0.00$.01) $0.31\pm 0.00$.02) $0.42\pm 0.00$.01) $0.31\pm 0.00$.01) Inc-3 $03\pm 0.00.00$) $0.13\pm 0.00$.01) $0.43\pm 0.00$.02) $0.52\pm 0.00$.02) $41\pm 0.00.00$) Mix-1 $03\pm 0.00.00$) $14\pm 0.00.00$) $0.49\pm 0.00$.01) $0.58\pm 0.00$.02) $0.51\pm 0.00$.02) Mix-2 $03\pm 0.00.00$) $0.13\pm 0.00$.01) $0.42\pm 0.00$.01) $0.50\pm 0.00$.02) $45\pm 0.00.00$) HOPSE-G Adj-1 $02\pm 0.00.00$) $0.09\pm 0.00$.01) $0.25\pm 0.00$.01) $0.26\pm 0.00$.01) $24\pm 0.00.00$) Adj-2 $02\pm 0.00.00$) $09\pm 0.00.00$) $0.29\pm 0.00$.01) $0.27\pm 0.00$.01) $0.26\pm 0.00$.01) Adj-3 $02\pm 0.00.00$) $10\pm 0.00.00$) $0.32\pm 0.00$.01) $0.32\pm 0.00$.02) $0.27\pm 0.00$.01) Inc-1 $02\pm 0.00.00$) $11\pm 0.00.00$) $0.37\pm 0.00$.01) $34\pm 0.00.00$) $33\pm 0.00.00$) Inc-2 $02\pm 0.00.00$) $09\pm 0.00.00$) $0.32\pm 0.00$.01) $0.30\pm 0.00$.05) $0.28\pm 0.00$.02) Inc-3 $02\pm 0.00.00$) $0.14\pm 0.00$.01) $40\pm 0.00.00$) $37\pm 0.00.00$) $0.36\pm 0.00$.02) Mix-1 $03\pm 0.00.00$) $14\pm 0.00.00$) $0.44\pm 0.00$.01) $45\pm 0.00.00$) $42\pm 0.00.00$) Mix-2 $02\pm 0.00.00$) $14\pm 0.00.00$) $0.40\pm 0.00$.02) $0.37\pm 0.00$.01) $0.32\pm 0.00$.01)
I.3 Preprocessing time
Tables˜13 and 12 show the preprocessing times for both HOPSE-M and HOPSE-G across various neighborhoods. While the tables indicate that HOPSE-M can be faster on these specific benchmarks, this empirical result requires careful contextualization and does not reflect the models’ true asymptotic scalability. The primary limitation of HOPSE-M is the high asymptotic complexity of its handcrafted encodings. Certain methods require operations like eigendecomposition, which scales at $O(n^{3})$ and becomes computationally prohibitive on larger graphs. In contrast, HOPSE-G is designed for greater scalability. Its preprocessing consists of a single GNN forward pass with more favorable scaling properties. Furthermore, HOPSE-G enhances downstream efficiency by consolidating all structural information into a single channel ($K=1$), which reduces the number of learnable projection functions ($f_{r,k}$) and leads to faster training and inference. Therefore, when considering both theoretical complexity and end-to-end efficiency, HOPSE-G is the more scalable framework. As expected, the tables show that preprocessing times increase with the complexity of the neighborhood definition, with the simple graph-based Adj-1 neighborhood being the fastest.
Table 12: Preprocessing time (seconds) of $\mathcal{N}_{S}$ on simplicial domain.
Graph Tasks Topological Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $(\beta_{0},\beta_{1},\beta_{2})$ HOPSE-M Adj-1 2.42 35.15 53.37 53.82 142.01 333.65 333.65 333.65 Adj-2 3.24 49.49 72.94 72.80 195.71 823.68 823.68 823.68 Adj-3 3.26 58.49 74.16 73.40 199.37 973.49 973.49 973.49 Inc-1 3.62 79.37 92.40 92.20 229.92 1121.78 1121.78 1121.78 Inc-2 4.39 80.91 111.65 111.49 280.18 1544.42 1544.42 1544.42 Inc-3 5.58 123.97 147.31 148.10 367.95 2233.27 2233.27 2233.27 Mix-1 7.27 165.71 191.94 193.45 489.55 3724.43 3724.43 3724.43 Mix-2 4.99 95.34 118.46 117.35 318.12 2495.14 2495.14 2495.14 HOPSE-G Adj-1 4.54 45.94 103.91 103.47 297.61 913.38 913.38 913.38 Adj-2 7.81 65.54 172.20 172.97 496.73 1676.79 1676.79 1676.79 Adj-3 7.76 92.15 174.23 174.49 502.85 2742.26 2742.26 2742.26 Inc-1 8.05 93.32 181.82 183.44 532.93 2654.36 2654.36 2654.36 Inc-2 11.21 90.87 245.87 248.52 714.05 2813.64 2813.64 2813.64 Inc-3 14.53 142.22 324.62 326.91 952.95 4432.76 4432.76 4432.76 Mix-1 21.02 210.41 474.58 472.70 1381.15 7329.51 7329.51 7329.51 Mix-2 14.28 160.55 314.86 319.78 932.91 5241.30 5241.30 5241.30
Table 13: Preprocessing time (seconds) of $\mathcal{N}_{S}$ on cell domain.
Graph Tasks $\mathcal{N}_{S}$ MUTAG PROTEINS NCI1 NCI109 ZINC HOPSE-M Adj-1 2.81 27.16 59.68 60.07 163.66 Adj-2 3.75 40.70 81.96 82.60 221.18 Adj-3 4.44 50.97 96.94 97.38 267.01 Inc-1 4.97 74.99 120.83 121.33 310.83 Inc-2 4.98 75.22 121.62 122.38 312.12 Inc-3 7.07 116.00 182.49 179.43 461.41 Mix-1 9.66 162.30 242.08 244.40 627.80 Mix-2 7.04 89.45 159.44 160.62 441.84 HOPSE-G Adj-1 5.08 38.04 113.21 108.81 315.07 Adj-2 8.34 58.50 189.71 184.54 525.01 Adj-3 11.10 75.56 253.63 248.51 727.66 Inc-1 11.89 85.26 264.02 259.75 734.69 Inc-2 11.81 120.32 292.64 257.43 740.39 Inc-3 18.51 132.25 420.45 403.67 1163.78 Mix-1 28.02 188.99 641.07 621.33 1776.32 Mix-2 20.47 140.60 472.04 460.79 1326.40
Appendix J Ablations over choices of PSEs
This section introduces an ablation study over the choices of PSEs as ways to characterize neighborhoods. The conditions include experiments for each individual PSE, a mix of all PSEs and a mix of all PSE without LapPE. The following results show these combinations over HOPSE-M and graph-based models GCN, GAT and GIN.
J.1 Over HOPSE-M
Table 14: Ablation in HOPSE-M using different PSEs on Graph Tasks.
PSE MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) Simplicial EStat $85.11\pm 0.06$.20) $74.70\pm 0.03$.03) $75.21\pm 0.01$.30) $74.04\pm 0.01$.59) $0.53\pm 0.00$.03) HKDiag $80.43\pm 0.04$.61) $76.13\pm 0.03$.09) $75.66\pm 0.00$.32) $75.90\pm 0.00$.48) $0.42\pm 0.00$.06) LapPE $78.30\pm 0.04$.09) $75.41\pm 0.01$.75) $73.99\pm 0.00$.90) $72.99\pm 0.00$.67) $64\pm 0.00.00$) RWSE $83.40\pm 0.07$.12) $75.20\pm 0.03$.48) $76.44\pm 0.00$.93) $74.39\pm 0.00$.86) $0.42\pm 0.00$.01) All w/o LapPE $86.38\pm 0.06$.31) $75.13\pm 0.03$.66) $77.35\pm 0.00$.81) $76.34\pm 0.01$.40) $0.39\pm 0.00$.02) Cell EStat $80.43\pm 0.05$.51) $75.20\pm 0.03$.12) $76.52\pm 0.02$.26) $76.19\pm 0.01$.75) $0.28\pm 0.00$.01) HKDiag $83.40\pm 0.06$.96) $76.99\pm 0.01$.94) $77.80\pm 0.00$.82) $76.53\pm 0.01$.58) $0.25\pm 0.00$.02) LapPE $58.94\pm 0.26$.35) $66.45\pm 0.09$.00) $62.95\pm 0.13$.52) $62.19\pm 0.12$.76) $35\pm 0.00.00$) RWSE $86.38\pm 0.05$.34) $73.33\pm 0.02$.34) $76.05\pm 0.01$.75) $75.45\pm 0.01$.45) $25\pm 0.00.00$) All w/o LapPE $88.09\pm 0.05$.55) $76.99\pm 0.02$.63) $78.09\pm 0.01$.44) $77.27\pm 0.01$.01) $0.25\pm 0.00$.01) All $85.53\pm 0.05$.51) $73.62\pm 0.02$.59) $76.32\pm 0.01$.74) $77.08\pm 0.01$.35) $0.24\pm 0.00$.01)
Table 15: Ablation in HOPSE-M using different PSEs on Topological Tasks.
PSE NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}(\uparrow)$ $\beta_{2}(\uparrow)$ EStat $81.69\pm 0.06$.38) $66.34\pm 0.12$.46) $90.15\pm 0.00$.30) $67.94\pm 0.01$.24) HKDiag $79.18\pm 0.06$.13) $77.27\pm 0.00$.91) $89.65\pm 0.00$.51) $59.59\pm 0.07$.16) LapPE $95.51\pm 0.00$.23) $89.23\pm 0.00$.15) $91.02\pm 0.01$.43) $81.47\pm 0.01$.16) RWSE $86.78\pm 0.07$.97) $78.59\pm 0.02$.46) $90.12\pm 0.00$.17) $65.81\pm 0.01$.40) All $96.27\pm 0.00$.28) $89.24\pm 0.00$.22) $90.87\pm 0.01$.07) $82.29\pm 0.00$.46)
J.2 Over GNNs
Table 16: Ablation on GNNs using different PSEs on Graph Tasks.
PSE MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) ZINC ($\downarrow$) GCN EStat $78.72\pm 0.04$.76) $74.84\pm 0.02$.96) $73.95\pm 0.01$.30) $73.51\pm 0.01$.32) $0.60\pm 0.00$.01) HKDiag $70.21\pm 0.05$.42) $74.27\pm 0.00$.64) $74.32\pm 0.00$.63) $72.26\pm 0.01$.27) $0.42\pm 0.00$.04) LapPE $60.99\pm 0.25$.49) $66.19\pm 0.07$.71) $62.73\pm 0.13$.28) $61.52\pm 0.12$.02) $0.62\pm 0.00$.01) RWSE $79.15\pm 0.09$.44) $73.62\pm 0.03$.56) $74.88\pm 0.00$.54) $74.70\pm 0.00$.79) $0.43\pm 0.00$.03) All w/o LapPE $77.02\pm 0.04$.09) $75.77\pm 0.02$.98) $73.42\pm 0.01$.37) $73.61\pm 0.01$.09) $0.48\pm 0.00$.03) All $84.26\pm 0.03$.56) $67.50\pm 0.08$.87) $62.03\pm 0.09$.91) $62.91\pm 0.10$.78) $0.58\pm 0.00$.05) GAT EStat $74.89\pm 0.04$.85) $74.84\pm 0.02$.02) $76.19\pm 0.01$.55) $74.48\pm 0.01$.05) $0.59\pm 0.00$.01) HKDiag $73.19\pm 0.04$.15) $74.70\pm 0.01$.55) $77.86\pm 0.01$.02) $75.57\pm 0.00$.60) $0.39\pm 0.00$.02) LapPE $61.23\pm 0.26$.12) $65.84\pm 0.08$.32) $62.55\pm 0.13$.11) $61.55\pm 0.12$.09) $0.61\pm 0.00$.01) RWSE $76.17\pm 0.08$.82) $73.84\pm 0.03$.16) $75.80\pm 0.01$.26) $73.57\pm 0.01$.07) $0.44\pm 0.00$.04) All w/o LapPE $78.30\pm 0.04$.61) $75.48\pm 0.03$.44) $75.76\pm 0.01$.68) $75.57\pm 0.01$.48) $0.47\pm 0.00$.04) All $80.43\pm 0.05$.08) $66.47\pm 0.08$.19) $62.91\pm 0.13$.48) $62.74\pm 0.13$.31) $0.56\pm 0.00$.02) GIN EStat $77.45\pm 0.08$.85) $73.91\pm 0.03$.31) $75.64\pm 0.01$.51) $74.81\pm 0.01$.00) $0.55\pm 0.00$.05) HKDiag $77.45\pm 0.10$.27) $74.77\pm 0.02$.67) $74.86\pm 0.00$.82) $74.39\pm 0.01$.11) $0.36\pm 0.00$.04) LapPE $60.99\pm 0.25$.69) $65.91\pm 0.08$.38) $62.63\pm 0.13$.17) $62.12\pm 0.12$.65) $0.53\pm 0.00$.02) RWSE $74.89\pm 0.10$.68) $72.83\pm 0.03$.23) $73.52\pm 0.02$.87) $74.00\pm 0.01$.69) $0.41\pm 0.00$.14) All w/o LapPE $76.17\pm 0.07$.12) $75.13\pm 0.02$.22) $74.30\pm 0.00$.84) $75.55\pm 0.01$.07) $0.44\pm 0.00$.02) All $83.83\pm 0.03$.56) $73.76\pm 0.03$.38) $75.84\pm 0.01$.51) $73.90\pm 0.01$.59) $0.54\pm 0.00$.10)
Table 17: Ablation on GNNs using different PSEs on Topological Tasks.
PSE NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}(\uparrow)$ $\beta_{2}(\uparrow)$ GIN EStat $81.39\pm 0.00$.92) $56.73\pm 0.01$.34) $88.27\pm 0.00$.10) $24.27\pm 0.04$.41) HKDiag $79.73\pm 0.01$.72) $56.27\pm 0.00$.45) $88.39\pm 0.00$.16) $27.50\pm 0.06$.52) LapPE $91.44\pm 0.01$.66) $81.68\pm 0.00$.96) $89.77\pm 0.00$.05) $63.65\pm 0.00$.49) RWSE $77.37\pm 0.02$.15) $68.89\pm 0.01$.48) $88.50\pm 0.00$.24) $32.82\pm 0.09$.71) All w/o LapPE $82.03\pm 0.00$.68) $60.09\pm 0.04$.87) $88.26\pm 0.00$.09) $24.20\pm 0.04$.31) All $94.05\pm 0.00$.15) $82.53\pm 0.00$.78) $90.28\pm 0.00$.05) $68.43\pm 0.00$.59)
Appendix K Model parameters.
Table 18: Model sizes (in thousands of parameters) corresponding to the best-performing hyperparameter configurations.
Model MUTAG PROTEINS NCI1 NCI109 ZINC NAME ORIENT $\beta_{0},\beta_{1},\beta_{2}$ Graph GCN 67.97K 13.19K 6.72K 23.75K 22.59K 3.56K 3.46K 3.49K GAT 22.02K 10.11K 11.20K 11.23K 22.85K 3.62K 3.52K 51.08K GIN 38.40K 13.19K 154.37K 154.50K 10.40K 3.56K 13.06K 13.12K GIN + PSEs 43.97K 29.82K 13.89K 16.22K 905.21K 128.26K 616.70K 485.37K GAT + PSEs 35.77K 13.15K 11.68K 176.38K 130.56K - - - GCN + PSEs 15.33K 12.03K 16.29K 33.47K 140.54K - - - Simplicial SCN 20.03K 10.24K 94.98K 26.08K 24.42K 348.55K 22.53K 348.29K SCCNN 27.11K 26.72K 188.99K 49.54K 1453.82K 362.88K 362.69K 1446.40K SaNN 286.08K 1523.97K 626.95K 632.19K 995.59K 891.14K 890.75K 890.88K GCCN 15.11K 46.27K 68.99K 49.63K 67.42K 1758.60K 953.35K 6957.31K HOPSE-M 1169.67K 1175.81K 2983.94K 809.86K 764.80K 579.46K 1198.08K 4593.92K HOPSE-G 359.04K 365.83K 1397.25K 1795.84K 666.50K 1370.37K 658.95K 2610.95K Cell CCCN 284.29K 34.56K 63.87K 17.67K 287.74K - - - CWN 334.72K 101.12K 124.10K 412.29K 88.06K - - - GCCN 15.11K 46.27K 68.99K 49.63K 39.78K - - - HOPSE-M 558.34K 2957.83K 809.47K 2189.06K 864.64K - - - HOPSE-G 1405.95K 357.89K 1842.69K 1445.63K 690.30K - - -
To ensure a fair comparison, our hyperparameter search, guided by the TopoBench framework, constrained all models to a comparable maximum parameter budget. The final parameter s for the best-performing configuration of each model are detailed in Table˜18. Interestingly, the search process often selected more compact models for the graph-based baselines, a finding consistent with the original TopoBench study. In contrast, the best-performing models for HOPSE were comparatively larger. We assume this is because HOPSE’s architecture is not constrained by the over-smoothing and over-squashing issues that typically limit the depth of message-passing-based methods. Despite its empirically larger size, HOPSE maintains a has a very competitive execution runtime, highlighting the scalability of its design. We believe this efficiency is a crucial feature for enabling the application of Topological Deep Learning on large and complex datasets.
Appendix L Licenses
This work does not distribute existing datasets. However, the dataset it uses will be mentioned here along with their licenses. The graph datasets MUTAG, PROTEINS, NCI1, NCI109 do not specify a license and are distributed as part of ’TUDatasets’ accessed via https://pubs.acs.org/doi/full/10.1021/acs.jcim.5b00559. The ZINC dataset from https://pubs.acs.org/doi/full/10.1021/acs.jcim.5b00559 is freely available. The MANTRA benchmarking dataset is licensed under the 3-Clause BSD license.
Appendix M Author contributions.
M.C. led the implementation, developed the theoretical framework, performed data analysis, and drafted the manuscript. M.M. primarily contributed to the theoretical framework and manuscript writing. N.M. provided infrastructure and project support, and contributed to writing. L.T. and G.B. conceived the research idea, designed the study, analyzed the results, and contributed to manuscript preparation. Additionally, L.T. participated in the implementation and theoretical development, conducted experiments with M.C., and supervised the project.